### v1

# Homological shifts of a complementary edge ideal

Abstract.
The homological shift algebra and the projective dimension function of complementary edge ideals are investigated. Let $G$ be a connected graph, and let $I$ be its complementary edge ideal. For bipartite graphs $G$, we show that the projective dimension of $I^{s}$ increases strictly with $s$ until reaching its maximum value. For trees and cycles, explicit expressions for the projective dimension of $I^{s}$ are provided, along with detailed descriptions of their homological shift algebras. In particular, it is shown that the $i$-th homological shift algebra of such ideals is generated in degree at most $i$. Additionally, we prove that if $G$ is a tree, then the homological shift ideal $\mathrm{HS}_{i}(I^{i})$, when divided by a suitable monomial, becomes a Veronese-type ideal, and every Veronese-type ideal arises in this manner.
2020 Mathematics Subject Classification. Primary 13A02, 13C15; Secondary 05E40, 05C05, 05C38
Keywords: projective dimension, tree, cycle, homological shift ideal, complementary edge ideal
Introduction
Let $S={\mathbb{K}}[x_{1},\ldots,x_{n}]$ be the polynomial ring in $n$ variables over a field $\mathbb{K}$}, and let $I\subset S$ be a non-zero, squarefree monomial ideal generated in degree $d$. Fröberg showed in [20] that if $d=2$, then $I$ has a linear resolution precisely when $I$ is the edge ideal of a chordal graph. Subsequently, Herzog, Hibi and Zheng proved in [9] that $I$ has a linear resolution if and only if all powers $I^{k}$ have linear quotients (also see [22, Theorem 3.2]). A monomial ideal $I$ is fully supported if for every $1\leq k\leq n$, the variable $x_{k}$ divides some minimal monomial generator of $I$. More recently, Ficarra and Moradi demonstrated in [15, Corollary 3.2] that if $I$ is fully supported in $S$ and is generated in degree d=n2d=n-2, then $I$ has a linear resolution if and only if $I^{k}$ has linear quotients for all $k\geq 1$. For any non-zero squarefree monomial ideal in degree n2n-2, it can be regarded as the complementary edge ideal of a finite simple graph $G$ with $n$ vertices. Let $G$ be a simple graph with with $n$ vertices, and let $E(G)$ be its edge set. We then define the complementary edge ideal of $G$ as
$I_{c}(G)=\left(\frac{{\mathbf{x}}_{[n]}}{x_{i}x_{j}}:\ \{i,j\}\in E(G)\right),$
where ${\mathbf{x}}_{[n]}=\prod\limits_{i=1}^{n}x_{i}$ and $n]=\{1,\ldots,n\}$. Furthermore, Ficarra and Moradi also showed in [15, Corollary 3.2] that $I_{c}(G)$ has a linear resolution if and only if $G$ is connected.
The complementary edge ideal of a graph was independently studied in 2025 in [16] and [28]. Hibi et al., as well as Ficarra and Moradi independently characterized sequentially Cohen-Macaulay, Cohen-Macaulay and Gorenstein complementary edge ideals. Furthermore, Ficarra and Moradi also characterized nearly Gorenstein and matroidal complementary edge ideals. They also computed the regularity and the asymptotic depth of powers of complementary edge ideals, as well as the index of their depth stability.
To study the syzygies and resolutions of monomial ideals, we use the homological shift ideal. This concept first appeared implicitly in Miller and Sturmfels’s 2005 book [30, Theorem 2.18] and was formally defined by Herzog et al. in 2020 [25].
Let ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$ and ${\mathbf{x}}^{{\mathbf{a}}}=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$. Given a monomial ideal $I\subset S$, the $i$-th homological shift ideals of $I$ is defined as for all $i\geq 0$:
$\operatorname{HS}_{i}(I)=({\mathbf{x}}^{{\mathbf{a}}}\mid\operatorname{Tor}_{i}^{S}({\mathbb{K}},I)_{{\mathbf{a}}}\neq 0).$
Note that $\mbox{HS}_{0}(I)=I$. Homological shift ideals encode critical multigraded data arising from the syzygies of $I$, see [2, 3, 10, 11, 12, 13, 14, 17, 18, 19, 25, 26, 31, 33]. A major motivation for studying homological shift ideals stems from the Bandari-Bayati-Herzog conjecture that being polymatroidal is a homological shift property, which has been answered positively in [6] recently. Inspired by conjectures from [2] and [25], Ficarra and Qureshi introduced the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I))$ in [19]. By definition, this algebra is given by
$\mathrm{HS}_{i}(\mathcal{R}(I))=\bigoplus_{s\geq 0}\mathrm{HS}_{i}(I^{s}).$
When $I$ has linear powers, they showed that $I\cdot\mathrm{HS}_{i}(I^{s})\subseteq\mathrm{HS}_{i}(I^{s+1})$ for all $s\geq 0$, which implies that $\mathrm{HS}_{i}(\mathcal{R}(I))$ is a graded $\mathcal{R}(I)$-module. Furthermore, it is finitely generated (see [19, Theorem 1.4]). The generating degree of homological shift algebras for a polymatroidal ideal has been investigated in [13]. Ficarra and Lu conjecture that, if $I$ is a polymatroidal ideal, then $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degrees $\leq i$ as an $\mathcal{R}(I)$-module for all $i>0$ (see [13, Conjecture 5.1]). They proved that this conjecture is true for special values of $i$ such as $i=1$, as well as for special polymatroidal ideals, including principal Borel ideals and polymatroidal ideals that satisfy the strong exchange property. Matroidal edge ideals are another example (see [5, 13]).
In this paper, we focus on the algebraic and homological properties of the complementary edge ideal of a graph and its powers. Specifically, we focus on the homological shift ideal and the projective dimension of powers of the complementary edge ideals of trees and cycles. Let $\mathrm{pd}(-)$ denote the projective dimension of an ideal -. Our results are as follows:
Theorem 0.1.
Let $G$ be a simple graph and let $I_{c}(G)$ be its complementary edge ideals.
  • (1)
    If $G$ is a tree with $n$ vertices, then
    $\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}s,&\hbox{ \text{if\ } $s\in[n-2]$;}\\ n-2,&\hbox{ \text{if\ } $s\geq n-2$.}\end{array}\right$.
  • (2)
    If $G$ is an even cycle with $2n$ vertices, then
    $\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}2s,&\hbox{\text{ if\ } $s\in[n-1]$;}\\ 2n-2,&\hbox{\text{ if\ } $s\geq n-1$.}\end{array}\right$.
  • (3)
    If $G$ is an odd cycle with $2n+1$ vertices, then
    $\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}2s,&\hbox{\text{ if\ } $s\in[n]$;}\\ 2n,&\hbox{\text{ if\ } $s>n$.}\end{array}\right$.
Proposition 0.2.
Let $G$ be a connected bipartite graph on vertex set $[n]$. Then $\mathrm{pd}(I_{c}(G)^{s})$ strictly increases until it reaches the maximal value n2n-2.
Let $\alpha$ denote the monomial $\prod\limits_{i=1}^{n}x_{i}$.
Theorem 0.3.
Let $G$ be a connected graph and let $I=I_{c}(G)$ be its complementary edge ideal. Then for all $k\geq 0$, we have
$\mathrm{HS}_{1}(I^{k+1})=I^{k}\cdot\mathrm{HS}_{1}(I).$
Theorem 0.4.
Let $G$ be a tree with $n$ vertices, labelled as in Remark 3.5, and let $I=I_{c}(G)$ be its complementary edge ideal. Define a map $\phi:[n-1]\rightarrow[n]$ such that $\phi(i)$ is the unique $j\in[n]$, with $j>i$ for which $x_{i}x_{j}\in E(G)$. Then, for any $1\leq i\leq s$, we have
$\mathrm{HS}_{i}(I^{s})=I^{s-i}\cdot\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}:\ F\subseteq[n-2]\text{\ and\ }|F|=i\right).$
As an immediate consequence of this result, if $G$ is tree, then the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I_{c}(G)))$ is generated in degree $i$ for all $1\leq i\leq n-2$.
Theorem 0.5.
Let $G$ be a tree let $I=I_{c}(G)$ be its complementary edge ideal. Then $\operatorname{HS}_{i}(I_{c}(G)^{i})$, when divided by a suitable monomial, is an ideal of Veronese type. Furthermore. every ideal of Veronese type arises in this manner.
Studying the homological shift algebra of the complementary edge ideal of a cycle is much more complicated. Regarding this topic, our results are as follows:
Theorem 0.6.
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. Then, for all integers $1\leq i<n$ and $s\geq\left\lfloor\frac{i}{2}\right\rfloor$ , the following equality holds:
$\mathrm{HS}_{i}(I^{s})=\sum_{\begin{subarray}{c}\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor\end{subarray}}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\right)$
Theorem 0.7.
Let $G$ be a cycle and let $I_{c}(G)$ be its complementary edge ideal, then $\mathrm{HS}_{i}(\mathcal{R}(I_{c}(G)))$ is generated in degree at most iqi-q for all $i>0$. Here, $q=\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}$.
The article is organized as follows: Section 1 provides the definitions and basic facts that will be used throughout the paper. In Section 2, we present a theorem on the homological shift ideal of independent significance. In Section 3, we compute the projective dimension of powers of the complementary edge ideals of trees and cycles. Furthermore, if $G$ is a connected bipartite graph, we prove that $\mathrm{pd}(I_{c}(G)^{s})$ strictly increases until it reaches its maximum value. In Section 4, we present the explicit expressions of homological shift ideals of powers for the complementary edge ideals of trees and cycles.
1. Preliminary
This section provides the definitions and basic facts used throughout the paper.
For a vector ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in{\mathbb{Z}}^{n}$, we set ${\mathbf{x}}^{{\mathbf{a}}}=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$. For a subset $F\subseteq[n]$, we set ${\mathbf{x}}_{F}=\prod_{i\in F}x_{i}$. Given positive integers $k,n$ with $k\leq n$, let $[k,n]$ denote the set $\{k,k+1,\ldots,n\}$. In particular, if $k=1$, $[k,n]$ is denoted by $[n]$.
Definition 1.1.
Let $I\subseteq R$ be a monomial ideal with minimal multigraded free $S$-resolution
$\mathbb{F}\ \ :\ \ 0\rightarrow F_{p}\rightarrow F_{p-1}\rightarrow\cdots\rightarrow F_{0}\rightarrow I\rightarrow 0,$
where $F_{i}=\bigoplus\limits_{j=1}^{\beta_{i}(I)}S(-{\mathbf{a}}_{i,j})$. The vectors ${\mathbf{a}}_{i,j}\in{\mathbb{Z}}^{n}$ are called the $i$th multigraded shifts of the resolution $\mathbb{F}$. The monomial ideal
$\mbox{HS}_{i}(I)=\ ({\mathbf{x}}^{{\mathbf{a}}_{i,j}}:\ j=1,\dots,\beta_{i}(I))$
is called the $i$th homological shift ideal of $I$.
Note that $\mbox{HS}_{0}(I)=I$ and $\mbox{HS}_{i}(I)=(0)$ for $i<0$ or $i>\operatorname{pd}(I)$.
Let $I$ be a monomial ideal, and let $\mathcal{G}(I)$ be its unique minimal set of monomial generators. We say that $I$ has linear quotients if there exists an order $u_{1},\ldots,u_{m}$ of $\mathcal{G}(I)$ such that the ideal $(u_{1},\ldots,u_{i-1}):(u_{i})$ is generated by variables, for all $i=2,\ldots,m$. By [29, Lemma 2.1], we can always assume that $\deg(u_{1})\leq\dots\leq\deg(u_{m})$. When $I$ has linear quotients, we set $\operatorname{set}(u_{1})=\emptyset$ and
$\operatorname{set}(u_{i})\ =\ \{\ell\ :\ x_{\ell}\in(u_{1},\dots,u_{i-1}):(u_{i})\},\quad\textup{for}\ i=2,\dots,m.$
By [27, Lemma 1.5], if $I\subset S$ has linear quotients, then
(1) $\operatorname{HS}_{i}(I)\ =\ ({\bf x}_{F}u:\ u\in\mathcal{G}(I),\ F\subset\operatorname{set}(u),\ |F|=i),$
where ${\bf x}_{F}=\prod_{j\in F}x_{j}$ if $F\neq\emptyset$, and ${\bf x}_{\emptyset}=1$ otherwise. From this, we can conclude that the projection dimension of $I$ is
(2) $\mathrm{pd}(I)\ =\max\{|\operatorname{set}(u)|:\ u\in\mathcal{G}(I)\}.$
Let $G$ be a finite simple graph with vertex set $V(G)=\{1,\ldots,n\}$ and edge set $E(G)$. The complementary edge ideal $I_{c}(G)$ of $G$ is defined as the following monomial ideals
$I_{c}(G)=\left(\frac{{\mathbf{x}}_{[n]}}{x_{i}x_{j}}:\ \{i,j\}\in E(G)\right).$
The following Lemma can be proved in the same way as [23, Lemma 10.3.5].
Lemma 1.2.
Let $I,J\subset S$ be monomial ideals with $d$-linear resolution and suppose that $J\subset I$. Then, for all ${\mathbf{a}}=(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$ and all $i\geq 0$, we have
$\beta_{i,{\mathbf{a}}}(J)\leq\beta_{i,{\mathbf{a}}}(I).$
In particular, for all $i\geq 0$, we have $\operatorname{HS}_{i}(J)\subseteq\operatorname{HS}_{i}(I)$.
This result has two consequences. For a graph $G$, a subgraph $H$ of $G$ is called a spanning subgraph if $V(H)=V(G)$.
Lemma 1.3.
Let $G$ be a connected graph and $H$ be a connected subgraph of $G$. Let $u=\prod_{i\in V(G)\setminus V(H)}x_{i}$. Then $u^{s}\cdot\operatorname{HS}_{i}(I_{c}(H)^{s})\subseteq\operatorname{HS}_{i}(I_{c}(G)^{s})$ for all $i,s\geq 1$. In particular, $\mathrm{pd}(I_{c}(H)^{s})\leq\mathrm{pd}(I_{c}(G)^{s})$.
Lemma 1.4.
Let $G$ be a connected graph. Then $\beta_{i}(I_{c}(G)^{s})\leq\beta_{i}(I_{c}(G)^{s+1})$ for all $s\geq 1$. In particular, the projective dimension function $\mathrm{pd}(I_{c}(G)^{s})$ of $I_{c}(G)$ is a monotonically increasing function.
Proof.
Fix a minimal monomial generator $u$ of $I_{c}(G)$. Then $uI_{c}(G)^{s}\subseteq I_{c}(G)^{s+1}$ and both have (n2)(s+1)(n-2)(s+1)-linear resolution. It follows that $\beta_{i}(I_{c}(G)^{s})=\beta_{i}(uI_{c}(G)^{s})\leq\beta_{i}(I_{c}(G)^{s+1})$ by Lemma 1.2.    ∎
2. A Theorem on the homological shift ideals of monomial ideals with linear resolutions
In this section, we present a result regarding the homological shift ideal of a monomial ideal with a linear resolution. Although this result is not directly related to other parts of this paper, it is interesting and applies to general monomial ideals. We therefore include it here.
Let $\mathfrak{m}$ be the maximal graded ideal of $S=\mathbb{K}[x_{1},\ldots,x_{n}]$, and let $\mathfrak{m}^{[i]}$ denote the $i$-th square-free power of $\mathfrak{m}$, i.e., $\mathfrak{m}^{[i]}=({\bf x}_{F}:\ F\subseteq[n],|F|=i).$
Theorem 2.1.
Let $I\subseteq S=\mathbb{K}[x_{1},\ldots,x_{n}]$ be a monomial ideal with a $d$-linear resolution. Then for all $i\in[n]$, we have
$\mathrm{HS}_{i}(I)+\mathrm{HS}_{i-1}(\mathfrak{m}I)=\mathfrak{m}^{[i]}I.$
Proof.
Consider the short exact sequence of multi-graded $S$-modules:
$0\rightarrow\mathfrak{m}I\rightarrow I\rightarrow T\rightarrow 0,$
where $T=\bigoplus\limits_{X^{\alpha}\in\mathcal{G}(I)}\mathbb{K}[-\alpha]$.
Fix an integer $i\in[n]$ and a multi-degree $\beta\in\mathbb{N}^{n}$ with $|\beta|=i+d$. Since $I$ has a $d$-linear resolution, its $k$-th syzygy module is concentrated in degree $d+k$ for all $k\geq 0$, so $\mathrm{Tor}_{i-1}(I,\mathbb{K})_{\beta}=0$. For $\mathfrak{m}I$, note that multiplying an ideal with a $d$-linear resolution by $\mathfrak{m}$ yields an ideal with a $(d+1)$-linear resolution, so $\mathrm{Tor}_{i}(\mathfrak{m}I,\mathbb{K})_{\beta}=0$.
Applying the $\mathbb{K}\otimes_{S}$- functor to the short exact sequence above, we obtain the following long exact sequence in multi-graded $\mathbb{K}$-vector spaces (truncated at the relevant degrees):
$\rightarrow\mathrm{Tor}_{i}(\mathfrak{m}I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i}(I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i}(T,\mathbb{K})_{\beta}\rightarrow\\ \mathrm{Tor}_{i-1}(\mathfrak{m}I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i-1}(I,\mathbb{K})_{\beta}\rightarrow.$
Since $\mathrm{Tor}_{i-1}(I,\mathbb{K})_{\beta}=\mathrm{Tor}_{i}(\mathfrak{m}I,\mathbb{K})_{\beta}=0$, this reduces to a short exact sequence:
$0\rightarrow\mathrm{Tor}_{i}(I,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i}(T,\mathbb{K})_{\beta}\rightarrow\mathrm{Tor}_{i-1}(\mathfrak{m}I,\mathbb{K})_{\beta}\rightarrow 0.$
Thus, $\mathrm{Tor}_{i}(T,\mathbb{K})_{\beta}\neq 0$ if and only if either $\mathrm{Tor}_{i}(I,\mathbb{K})_{\beta}\neq 0$ or $\mathrm{Tor}_{i-1}(\mathfrak{m}I,\mathbb{K})_{\beta}\neq 0$. Since all minimal generators of $\mathrm{HS}_{i}(I)$, $\mathrm{HS}_{i-1}(\mathfrak{m}I)$, and $\mathrm{HS}_{i}(T)$ have degree $i+d$ (matching $|\beta|$), we conclude that
$\mathrm{HS}_{i}(I)+\mathrm{HS}_{i-1}(\mathfrak{m}I)=\mathrm{HS}_{i}(T).$
On the other hand, the Koszul complex of $x_{1},\ldots,x_{n}$ provides a minimal free resolution of $\mathbb{K}$. It follows that for each $X^{\alpha}\in\mathcal{G}(I)$, $\mathrm{Tor}_{i}(\mathbb{K}[-\alpha],\mathbb{K})_{\beta}\neq 0$ if and only if $\beta=\alpha+\gamma$ for some $\gamma\in\mathbb{N}^{n}$ with $X^{\gamma}\in\mathcal{G}(\mathfrak{m}^{[i]})$. Since $T=\bigoplus_{X^{\alpha}\in\mathcal{G}(I)}\mathbb{K}[-\alpha]$, $\mathrm{HS}_{i}(T)$ is generated by all monomials $X^{\alpha+\gamma}=X^{\alpha}X^{\gamma}$ where $X^{\alpha}\in\mathcal{G}(I)$ and $X^{\gamma}\in\mathcal{G}(\mathfrak{m}^{[i]})$, which is exactly $\mathfrak{m}^{[i]}I$.
Combining the two equalities, we get $\mathrm{HS}_{i}(I)+\mathrm{HS}_{i-1}(\mathfrak{m}I)=\mathfrak{m}^{[i]}I$.    ∎
This result indicates that every minimal generator of $\mathrm{HS}_{i}(I)$ has the form ${\bf x}_{F}\cdot u$, where $u\in\mathcal{G}(I)$ and $F$ is a subset of cardinality $i$ in $[n]$. This is consistent with the case when $I$ has linear quotients. In [14], the authors expected that if $I$ has linear resolutions then $\mathrm{HS}_{1}(I)$ would have linear quotients (see [14, the third paragraph, p134]). This result lends support to this expectation.
Another consequence of Theorem 2.1 is as follows:
Corollary 2.2.
Let $I\subset S$ be a monomial ideal with a $d$-linear resolution. Then $\mathrm{HS}_{n-1}(\mathfrak{m}I)=\mathfrak{m}^{[n]}I$. In particular, if $I\neq 0$, then $\mathrm{pd}(\mathfrak{m}I)=n-1$.
Proof.
Note that $\mathrm{HS}_{n}(I)=0$. Specializing $i$ to $n$ in Theorem 2.1 yields the desired formula.    ∎
3. Projective dimensions functions
In this section, we determine the projective dimension of powers of complementary edge ideals for several classes of connected graphs, including trees and cycles.
First, we will review the concept of even-connectedness and a related result. By abuse of notation, an edge of $G$ refers to both a two-element subset $\{i,j\}$ of $V(G)$ and a monomial $x_{i}x_{j}$. This abuse of notation should not cause any confusion.
Definition 3.1.
([1, Definition 6.2]) Two vertices $j$ and $k$ ($j$ may be equal to $k$) of $G$ are said to be even-connected with respect to the $s$-fold product $e_{1}\cdots e_{s}$ if there exists a walk $j_{1}--j_{2}--\cdots--j_{2\ell+2}$ ($\ell\geq 2$) in $G$ such that
  1. (1)
    $j_{1}=j$ and $j_{2\ell+2}=k$;
  2. (2)
    For all $t\in[\ell]$, $j_{2t}j_{2t+1}=e_{i}$ for some $i$;
  3. (3)
    For all $i\in[s]$, $|\{t\in[\ell]:\ j_{2t}j_{2t+1}=e_{i}\}|\leq|\{t\in[s]:\ e_{t}=e_{i}\}|$.
Theorem 3.2.
([1, Theorems 6.1 and 6.7]) Let $G$ be a graph and let $e_{1}\cdots e_{s}$ be an $s$-fold product of edges in $G$, where $s\geq 1$ and $e_{i}$ may be equal to $e_{j}$ for $i\neq j$. Then $(I(G)^{s+1}:e_{1}\cdots e_{s})$ is generated by monomials of degree two. Moreover, each generator $x_{j}x_{k}$ ($x_{j}$ may be equal to $x_{k}$) of $(I(G)^{s+1}:e_{1}\cdots e_{s})$ is either an edge of $G$ or even-connected with respect to $e_{1}\cdots e_{s}$.
Let $G$ be a connected graph with vertex set $[n]$. As noted in [15, Remark 3.3], if the vertices of $G$ are labeled with integers from $[n]$ such that for all $i\in[n-1]$, the induced subgraph of $G$ on $V(G)\setminus[i]$ (i.e., on vertex set $\{i+1,\ldots,n\}$) is connected, then $I_{c}(G)^{k}$ has linear quotients for all $k\geq 1$ with respect to the lexicographical order on the minimal generators of $I_{c}(G)^{k}$ induced by $x_{1}>x_{2}>\cdots>x_{n}$. For the rest of this section, any order involved refers to this lexicographical order.
The following result plays a key role in this section and next section. For convenience, we always denote $\prod_{i\in V(G)}x_{i}$ by $\alpha(G)$ or simply $\alpha$.
Lemma 3.3.
Let $\beta\in\mathcal{G}(I_{c}(G)^{s})$, and write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$, where each $u_{i}\in\mathcal{G}(I(G))$, and $I(G)$ is the edge ideal of $G$. Then
$\mathrm{set}(\beta)\subseteq\{i\mid x_{i}\mbox{ divides }u_{1}\cdots u_{s}\}.$
Furthermore, if $u_{1}=x_{i}x_{j}$, then $i\in\mathrm{set}(\beta)$ if and only if there is some $k\in[n]$ with $k>i$ such that either $\{j,k\}\in E(G)$ or $j$ is even-connected to $k$ with respect to the (s1)(s-1)-fold product $u_{2}\cdots u_{s}$.
Proof.
For any monomials $U_{1},\ldots,U_{k},U$, one has $(\frac{\alpha^{s}}{U_{1}},\cdots,\frac{\alpha^{s}}{U_{k}}):\frac{\alpha^{s}}{U}=(U:U_{1})+\cdots+(U:U_{k})$. The first assertion follows from this observation. Next, we prove the second assertion.
Let $u=u_{1}\cdots u_{s}$. If $i\in\mathrm{set}(\beta)$, then there is some $\gamma>\beta$ such that $\gamma=\alpha^{s}/v$ and $\gamma:\beta=u:v=(x_{i})$, where $v\in\mathcal{G}(I(G)^{s})$. Note that $u=x_{i}x_{j}u_{2}\cdots u_{s}$, we can deduce that $v=x_{k}x_{j}u_{2}\cdots u_{s}$ for some $k\in[n]$ and that $u>v$, which implies that $i<k$. It is clear that $x_{k}x_{j}\in(I(G)^{s}:u_{2}\cdots u_{s})$, so by Theorem 3.2, either $\{j,k\}\in E(G)$ or $j$ is even-connected to $k$ with respect to the (s1)(s-1)-fold product $u_{2}\cdots u_{s}$.
Conversely, if there is $k\in[n]$ such that $k>i$, and either $\{j,k\}\in E(G)$ or $j$ is even-connected to $k$ with respect to the (s1)(s-1)-fold product $u_{2}\cdots u_{s}$, then $x_{k}x_{j}\in(I(G)^{s}:u_{2}\cdots u_{s})$. Let $v=x_{k}x_{j}u_{2}\cdots u_{s}$ and $\gamma=\alpha^{s}/v$. Then $\gamma>\beta$ and $\gamma\in\mathcal{G}(I_{c}(G)^{s})$. Clearly, $\gamma:\beta=(x_{i})$, so $i\in\mathrm{set}(\beta)$.    ∎
By [17, Propositions 4.3, 4.5], if $G$ is a connected graph, then the maximal value of $\mathrm{pd}(I_{c}(G)^{s})$ for $s\geq 1$ is n1n-1 if $G$ is non-bipartite, and n2n-2 if $G$ is bipartite. We first determine the projective dimension function of the complementary edge ideal of a cycle.
Theorem 3.4.
(1) Let $G$ be an even cycle with $2m$ vertices. Then
$\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}2s,&\hbox{if }s\in[m-1];\\ 2m-2,&\hbox{if }s\geq m-1.\end{array}\right$.
(2) Let $G$ be an odd cycle with $2m+1$ vertices. Then
$\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}2s,&\hbox{if }s\in[m];\\ 2m,&\hbox{if }s>m.\end{array}\right$.
Proof.
(1) Suppose the vertices of $G$ are labeled as $1-2-3-\cdots-(2m-1)-(2m)-1$. Fix an $s\in[m-1]$. For any $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}\in\mathcal{G}(I_{c}(G)^{s})$ where each $u_{i}\in\mathcal{G}(I(G))$, Lemma 3.3 implies $\mathrm{set}(\beta)\subseteq\{i\mid x_{i}\mbox{ divides }u_{1}u_{2}\cdots u_{s}\}$. Thus $|\mathrm{set}(\beta)|\leq 2s$, so $\mathrm{pd}(I_{c}(G)^{s})=\max\{|\mathrm{set}(\alpha)|:\alpha\in\mathcal{G}(I_{c}(G)^{s})\}\leq 2s$.
Conversely, we choose $u_{i}=x_{2i-1}x_{2i}$ for any $i\in[s]$, and set $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$. Then $\beta\in\mathcal{G}(I_{c}(G)^{s})$. By Lemma 3.3, $(2i-1)\in\mathrm{set}(\beta)$ for any $i\in[s]$. Note that $2m$ is even-connected to 2i12i-1 with respect to $u_{1}u_{2}\cdots u_{2i-3}$ via the walk:
$(2m)-1-2-3-4-\cdots-(2i-3)-(2i-2)-(2i-1).$
Given this fact and since $2m>2i$, we conclude $2i\in\mathrm{set}(\beta)$ by Lemma 3.3. Therefore, $\mathrm{set}(\beta)=[2s]$ and $\mathrm{pd}(I_{c}(G)^{s})=\max\{|\operatorname{set}(\alpha)|:\alpha\in\mathcal{G}(I_{c}(G)^{s})\}\geq 2s$.
Since $\mathrm{pd}(I_{c}(G)^{s})$ is increasing by Lemma 1.3, and $\mathrm{pd}(I_{c}(G)^{s})\leq 2m-2$ for all $s\geq 1$ by [17, Proposition 4.3], it follows that $\mathrm{pd}(I_{c}(G)^{s})=2m-2$ for all $s\geq m-1$.
The proof of (2) is similar to that of (1), so we omit the proof.    ∎
Remark 3.5.
Let $G$ be a tree with $n$ vertices. We fix a degree-one vertex of $G$ and label it $n$. There exists at least one labeling of the remaining vertices in $V(G)$ using the elements of [n1][n-1] such that $\text{dist}(i,n)\geq\text{dist}(j,n)$ for all $1\leq i<j\leq n-1$. Here, $\text{dist}(i,n)$ is the length of the shortest path between $i$ and $n$ (commonly referred to as “distance” in graph theory). In such a labeling, for each $i\in[n-1]$, there is a unique $k$ with $i<k\leq n$ such that $i$ is adjacent to $k$. This kind of labeling is not unique in general, and is unique precisely when $G$ is a path. In the rest of this section, all trees will be labeled in this manner.
Lemma 3.6.
Let $G$ be a tree and $\beta\in\mathcal{G}(I_{c}(G)^{s})$. Write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$, where each $u_{i}\in\mathcal{G}(I(G))$, Assume that $u_{1}=x_{i}x_{j}$ with $1\leq i<j\leq n$. Then
  1. (1)
    $i\in\mathrm{set}(\beta)$ if and only if $j\neq n$. In particular, $\mathrm{set}(\beta)\subseteq[n-2]$.
  2. (2)
    If $j\in[n-2]$, then $j\in\mathrm{set}(\beta)$ if and only if there is some $2\leq t\leq s$ such that $u_{t}=x_{j}x_{k}$ and $k>j$.
Proof.
(1) If $j\neq n$, then there is some $k\in[n]$ such that $k>j$ and $k$ is adjacent to $j$. Since $j\geq i$, we have $k>i$. By Lemma 3.3, $i\in\mathrm{set}(\beta)$.
If $j=n$, then i=n1i=n-1, since $n$ is a leaf of $G$. If $(n-1)\in\mathrm{set}(\beta)$, then by Lemma 3.3, $n$ is even-connected to $k$ with respect to the (s1)(s-1)-fold product $u_{2}\cdots u_{s}$. Thus there exists a closed walk $y_{1}-y_{2}-\cdots-y_{2\ell+1}-y_{2\ell+2}$ such that $y_{1}=y_{2\ell+2}=n$. However, since $G$ is a tree, every edge of $G$ appears an even number of times in this walk. Therefore, the walk is of even length. However, this walk has length $2\ell+1$, which is a contradiction. This implies $i\notin\mathrm{set}(\beta)$.
(2) If $j\in\mathrm{set}(\beta)$, then, by Lemma 3.3, there is some $k^{\prime}\in[n]$ such that $k^{\prime}>j$ and either $\{i,k^{\prime}\}\in E(G)$ or $i$ is even-connected to $k^{\prime}$ with respect to (s1)(s-1)-fold product $u_{2}\cdots u_{s}$. Since $G$ is a tree, based on the labeling of $G$’s vertices as described in Remark 3.5, we know that $j$ is the unique vertex satisfying $\{j,i\}\in E(G)$ and $j>i$. Thus $i$ is even-connected to $k^{\prime}$ with respect to (s1)(s-1)-fold product $u_{2}\cdots u_{s}$ by the walk $y_{1}-y_{2}-\cdots-y_{2\ell+1}-y_{2\ell+2}$. Let $y_{1}=i$ and $y_{2\ell+2}=k^{\prime}$, and let $p$ be the largest integer satisfying $y_{p}=i$. If $p\neq 1$, then $y_{1}-y_{2}-\cdots-y_{p-1}-y_{p}$ is a closed walk. This implies that $p$ is odd and $x_{y_{p+1}}x_{y_{p+2}}$ belongs to $\{u_{2},\ldots,u_{s}\}$. Again based on the labeling of $G$’s vertices as described in Remark 3.5, we know that $j=y_{p+1}$. If $y_{p+2}>j$, we can set $k=y_{p+2}$ and we are done. If $y_{p+2}<j$, let $q$ be the largest integer satisfying $y_{q}=j$ and $q\geq p+1$. Then q(p+1)q-(p+1) is even and $y_{q+1}>j$, since $G$ is a tree. We put $k=y_{p+1}$ and the result follows. The converse conclusion follows from (1).    ∎
Let $G$ be a graph and $e=\{i,j\}$ be an edge of $G$, we set $\min(e)=\min\{i,j\}$.
Theorem 3.7.
Let $G$ be a tree with $n$ vertices. Then
$\mathrm{pd}(I_{c}(G)^{s})=\left\{\begin{array}[]{ll}s,&\hbox{ \text{if\ } $s\in[n-2]$;}\\ n-2,&\hbox{ \text{if\ } $s\geq n-2$.}\end{array}\right$.
Proof.
Let $e_{1},e_{2},\ldots,e_{n-2},e_{n-1}$ be all the edges of $G$, where $e_{n-1}=\{n-1,n\}=x_{n-1}x_{n}$. Then $\min(e_{1}),\ldots,\min(e_{n-2})$ are pairwise distinct.
For any $s\in[n-2]$, we choose $\beta_{s}=\frac{\alpha^{s}}{e_{1}\cdots e_{s}}$. Then $\beta_{s}\in\mathcal{G}(I_{c}(G)^{s})$ and, by Lemma 3.6.(1), we can deduce that
$\{\min(e_{1}),\ldots,\min(e_{s})\}\subseteq\mathrm{set}(\beta_{s}).$
This implies that $\mathrm{pd}(I_{c}(G)^{s})\geq s$.
For any $\beta\in\mathcal{G}(I_{c}(G)^{s})$, we can write $\beta$ as $\beta=\frac{\alpha^{s}}{e_{i_{1}}\cdots e_{i_{s}}}$, where each $i_{j}\in[n-1]$. If $i\in\mathrm{set}(\beta)$, then by Lemma 3.6 (2), there is a $t\in[s]$ such that $i=\min(u_{t})$. This implies
(3) $\mathrm{set}(\beta)=\{\min(e_{i_{j}}))\mid j\in[s]\}\setminus\{n-1\}.$
Therefore, $|\mathrm{set}(\beta)|\leq s$. It follows from the arbitrariness of $\beta$ that $\mathrm{pd}(I_{c}(G)^{s})=\max\{|\operatorname{set}(\beta)|:\ \beta\in\mathcal{G}(I_{c}(G)^{s})\}\leq s$.
Suppose that $s\geq n-2$. Then
$n-2=\mathrm{pd}(I_{c}(G)^{n-2})\leq\mathrm{pd}(I_{c}(G)^{s})\leq n-2.$
This implies $\mathrm{pd}(I_{c}(G)^{s})=n-2$, as desired.    ∎
This result generalizes [17, Proposition 4.6], which proves that the same formula holds for paths, a specific type of trees.
The projective dimension of $I_{c}(G)$ for all connected graph is given.
Proposition 3.8.
Let $G$ be a connected graph. Then $\mathrm{pd}(I_{c}(G))$ is either $1$ or $2$. Moreover, $\mathrm{pd}(I_{c}(G))=1$ if and only if $G$ is a tree.
Proof.
It is straightforward to see that $1\leq\mathrm{pd}(I_{c}(G))\leq 2$ by the first assertion of Lemma 3.3.
If $G$ is a tree, then $\mathrm{pd}(I_{c}(G))=1$ by Theorem 3.7. If $G$ is not a tree, there exists a cycle $C$ that is a subgraph of $G$. By Lemma 3.4 and Lemma 1.3, we have $\mathrm{pd}(I_{c}(G))\geq\mathrm{pd}(I_{c}(C))=2$.    ∎
For any connected graph $G$, it is known that $\mathrm{pd}(I_{c}(G)^{s})$ is non-decreasing and $\mathrm{depth}(I_{c}(G)^{s})$ is non-increasing in $s$. This can be improved upon when $G$ is bipartite.
Proposition 3.9.
Let $G$ be a connected bipartite graph with $n$ vertices. Then $\mathrm{pd}(I_{c}(G)^{s})$ strictly increases until it reaches its maximum value n2n-2.
Proof.
By [17, Proposition 4.3] and the Auslander-Buchsbaum formula, we know that $\mathrm{pd}(I_{c}(G)^{s})=n-2$ for all $s\gg 0$. Suppose that $\mathrm{pd}(I_{c}(G)^{s})=k\leq n-3$. We can choose $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in\mathcal{G}(I_{c}(G)^{s})$ such that $|\operatorname{set}(\beta)|=k$. Let $\ell$ be the minimal element of $n]\setminus\operatorname{set}(\beta)$, and let $T$ be a spanning tree of $G$, and label its vertices as described in Remark 3.5. Since $k\leq n-3$, it follows that $\ell\leq n-2$. Thus, there exists an integer $m$ with $\ell<m\leq n-1$ such that $\{\ell,m\}\in E(T)$. Set $u_{s+1}=x_{\ell}x_{m}$ and $\gamma=\beta\cdot\frac{\alpha}{u_{s+1}}$. By Lemma 3.3, $\ell\in\operatorname{set}(\gamma)$. Therefore, $|\operatorname{set}(\gamma)|\geq k+1$. It follows that $\mathrm{pd}(I_{c}(G)^{s+1})\geq k+1$, which completes the proof.    ∎
If $G$ is a connected non-bipartite graph with $n$ vertices, then the maximal value of $\mathrm{pd}(I_{c}(G)^{s})$ is n1n-1. Using the same method as in Lemma 3.9, we can prove that $\mathrm{pd}(I_{c}(G)^{s})$ is strictly increasing in $s$ until it reaches n2n-2. However, it is difficult to prove that it continues to increase strictly when it reaches n2n-2.
By [4], we know that the limit $\lim\limits_{k\rightarrow\infty}\mathrm{depth}(S/I^{k})$ exists for any ideal $I\subset S$, that is, $\mathrm{depth}(S/I^{k})=\mathrm{depth}(S/I^{k+1})$ for all $k\gg 0$. The least integer $k_{0}>0$ for which $\mathrm{depth}(S/I^{k})=\mathrm{depth}(S/I^{k_{0}})$ for all $k\geq k_{0}$ is called the index of depth stability of $I$ and is denoted by $\mathrm{dstab}(I)$.
Corollary 3.10.
Let $G$ be a connected bipartite graph with $n$ vertices. Then
$\mathrm{dstab}(I_{c}(G))\leq n-2.$
The equality holds if and only if $G$ is a tree.
Proof.
By Theorem 3.7 together with the Auslander-Buchsbaum formula, we can show that if $G$ is a tree, then $\mathrm{dstab}(I_{c}(G))=n-2$.
Suppose that $G$ contains at least one cycle, denoted $C$. It suffices to show that $\mathrm{dstab}(I_{c}(G))<n-2$. By Lemma 3.4, $\mathrm{pd}(I_{c}(C))=2$. By Lemma 1.3, it follows that $\mathrm{pd}(I_{c}(G))\geq\mathrm{pd}(I_{c}(C))=2$. Assume, on the contrary, that $\mathrm{dstab}(I_{c}(G))\geq n-2$. Then, according to Proposition 3.9, we have
$\mathrm{pd}(I_{c}(G)^{n-2})\geq\mathrm{pd}(I_{c}(G)^{n-3})+1\geq\cdots\geq\mathrm{pd}(I_{c}(G))+(n-3)\geq 2+(n-3)=n-1.$
This yields a contradiction, completing the proof.    ∎
4. Homological shift algebras
In this section, we will examine the following conjecture and prove that it is true for graphs that are either trees or cycles.
Conjecture 4.1.
Let $G$ be a connected graph, and let $I=I_{c}(G)$ be its complementary edge ideal. Then $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degree $\leq i$.
4.1. The case when $i=1$
As proven in [13], for polymatroidal ideals $I$, $\mathrm{HS}_{1}(\mathcal{R}(I))$ is generated in degree $1$ with a relatively straightforward proof. The case $i=1$ for complementary edge ideals also admits a similarly elementary proof.
Theorem 4.2.
Let $G$ be a connected graph on the set $[n]$ and let $I=I_{c}(G)$ be its complementary edge ideal. Then $\mathrm{HS}_{1}(\mathcal{R}(I))$ is generated in degree $1$.
Proof.
It suffices to prove that $\mathrm{HS}_{1}(I^{s+1})=I^{s}\cdot\mathrm{HS}_{1}(I)$ for all $s\geq 0$. Let $\alpha=\prod\limits_{i=1}^{n}x_{i}$. It is easy to see that $\mathrm{HS}_{1}(I)=\left(\frac{\alpha}{x_{i}}\mid\deg_{G}(i)\geq 2\right)$. We will prove the statement by induction on $s$, and the case where $s=0$ is trivial. Suppose that $s\geq 1$. For any $\beta\in\mathcal{G}(\mathrm{HS}_{1}(I^{s+1}))$. Then, by [18, Lemma 2.3], there are two elements $u,v\in\mathcal{G}(I^{s+1})$ such that $\beta=\frac{\alpha^{s+1}}{\gcd(u,v)}$, where $\deg(\gcd(u,v))=2s-1$. We can write $u$ as $u=\prod\limits_{i=1}^{s+1}u_{i}$ where each $u_{i}\in\mathcal{G}(I(G))$. Therefore, $u=y_{1}\gcd(u,v)$. Without loss of generality, we can assume that $y_{1}$ divides $u_{1}$ and that $u_{1}=x_{i}y_{1}$ for some $i\in[n]$. Then $\gcd(u,v)=x_{i}\prod\limits_{i=2}^{s+1}u_{i}$. We consider the following two cases:
(1) If $\deg_{G}(i)\geq 2$, then $\beta$ can be written as
$\beta=\frac{\alpha^{s}}{u_{2}\cdots u_{s+1}}\cdot\frac{\alpha}{x_{1}},$
and the conclusion follows.
(2) If $\deg_{G}(i)=1$, then $v$ can be written as $v=\prod\limits_{i=1}^{s+1}v_{i}$. Assuming that $x_{i}$ divides $v_{1}$, we have $u_{1}=v_{1}$. In this case, we set $f=\prod\limits_{i=2}^{s+1}u_{i}$ and $g=\prod\limits_{i=2}^{s+1}v_{i}$. Then $\gcd(u,v)=u_{1}\gcd(f,g)$. Therefore, $\frac{\alpha^{s}}{\gcd(f,g)}\in\mathcal{G}(\mathrm{HS}_{1}(I^{s}))$. By the induction hypothesis, $\frac{\alpha^{s}}{\gcd(f,g)}\in\mathcal{G}(I^{s-1}\cdot\mathrm{HS}_{1}(I))$. Therefore, $\beta=\frac{\alpha}{u_{1}}\cdot\frac{\alpha^{s}}{\gcd(f,g)}\in\mathcal{G}(I^{s}\cdot\mathrm{HS}_{1}(I))$, as desired.    ∎
For $i\geq 2$, there is no uniform method to determine the generating degree of the $i$-th homological shift algebra of the complementary edge ideal of a connected graph.
4.2. Trees
Theorem 4.3.
Let $G$ be a tree with $n$ vertices, labelled as in Remark 3.5, and let $I=I_{c}(G)$ denote its complementary edge ideal. Define a map $\phi:[n-1]\rightarrow[n]$ such that $\phi(k)$ is the unique $j\in[n]$, with $j>k$ for which $\{k,j\}\in E(G)$. Then, for any $1\leq i\leq s$, we have
$\mathrm{HS}_{i}(I^{s})=I^{s-i}\cdot\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}:\ F\subseteq[n-2]\text{\ and\ }|F|=i\right).$
Proof.
Let $e_{1},e_{2},\ldots,e_{n-2},e_{n-1}$ be all the edges of $G$. Without loss of generality, we can assume that $\min(e_{i})=i$ for all $i\in[n-1]$. Thus, $e_{i}=x_{i}x_{\phi(i)}$. Note that every element of $\mathcal{G}(I(G)^{s})$ is uniquely written as
$e_{1}^{s_{1}}e_{2}^{s_{2}}\cdots e_{n-1}^{s_{n-1}},$
where $\sum\limits_{i=1}^{n-1}s_{i}=s$ and each $s_{i}\geq 0$. Furthermore, $e_{1},\ldots,e_{n-1}$ are algebraically independent.
For any $\beta=\frac{\alpha^{s-i}}{e_{1}^{a_{1}}\cdots e_{n-1}^{a_{n-1}}}\in\mathcal{G}(I^{s-i})$ and for any $F\subseteq[n-2]$ with $|F|=i$, we define
$\gamma=\frac{\alpha^{s}}{e_{1}^{a_{1}+b_{1}}\cdots e_{n-1}^{a_{n-1}+b_{n-1}}},$
where $b_{i}=1$ if $i\in F$, and $b_{i}=0$ otherwise. In particular, $b_{n-1}=0$, so $\gamma\in\mathcal{G}(I^{s})$, and $F\subseteq\operatorname{set}(\gamma)$ by the equation (3). This implies that ${\bf x}_{F}\gamma\in\mathrm{HS}_{i}(I^{s})$. Note that
${\bf x}_{F}\gamma=\beta\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}},$
we conclude that the inclusion $\supseteq$ holds.
Conversely, for any $U\in\mathcal{G}(\mathrm{HS}_{i}(I^{s}))$, we can write $U$ as
$U={\bf x}_{F}V={\bf x}_{F}\frac{\alpha^{s}}{e_{1}^{s_{1}}\cdots e_{n-1}^{s_{n-1}}},$
where $V=\frac{\alpha^{s}}{e_{1}^{s_{1}}\cdots e_{n-1}^{s_{n-1}}}\in\mathcal{G}(I^{s})$ and $F\subseteq\operatorname{set}(V)$ with $|F|=i$. By the equation (3), we see that $s_{i}>0$ for all $i\in F$. Let $c_{i}=s_{i}-1$ if $i\in F$, and $c_{i}=s_{i}$ otherwise. Then $W:=\frac{\alpha^{s-i}}{e_{1}^{c_{1}}\cdots e_{n-1}^{c_{n-1}}}\in\mathcal{G}(I^{s-i})$. Furthermore, we have $W\frac{\alpha^{i}}{\prod_{i\in F}x_{\phi(i)}}={\bf x}_{F}V$. This implies that the inclusion $\subseteq$ holds, completing the proof.    ∎
Corollary 4.4.
Let $G$ be a tree and let $I=I_{c}(G)$ be its complementary edge ideal. Then the $i$-th homological shift algebra $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degree $i$ for $1\leq i\leq n-2$.
Proof.
By Theorem 4.3, it follows that $\mathrm{HS}_{i}(I^{s})=I^{s-i}\cdot\mathrm{HS}_{i}(I^{i})$ for all $s\geq i$. This implies $\mathrm{HS}_{i}(\mathcal{R}(I))$ is generated in degree $\leq i$. But $\mathrm{HS}_{i}(I^{s})=0$ for $s<i$, the result follows.    ∎
The ideal $\left(\frac{\alpha^{i}}{\prod_{k\in F}x_{\phi(k)}}:\ F\subseteq[n-2]\text{\ and\ }|F|=i\right)$ is uniquely determined by $G$ and $i$ up to the permutation of variables, since it equals to $\mathrm{HS}_{i}(I^{i})$. This implies it is independent of the labelling of $G$. We denote this ideal by $J_{i}(G)$.
To study the properties of $J_{i}(G)$, we recall some theory on polymatroidal ideals. We follow the notation from [23] and [6]. Given a vector ${\bf a}=(a_{1},\dots,a_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$ and an integer $d$ with $a_{1}+\dots+a_{n}\geq d$, the ideal of Veronese type (determined by $\bf a$} and $d$) is the polymatroidal ideal $I_{{\bf a},d}$ with minimal monomial generating set
$\mathcal{G}(I_{{\bf a},d})\ =\ \{{\bf x^{b}}\in S:\ |{\bf b}|=d,\ {\bf b}\leq{\bf a}\}.$
Here, if ${\bf b}=(b_{1},\dots,b_{n})\in{\mathbb{Z}}_{\geq 0}^{n}$ we put ${\bf b}(i)=b_{i}$, $|{\bf b}|=\sum_{i=1}^{n}{\bf b}(i)$ is the modulus of $\bf b$} and ${\bf b}\leq{\bf a}$ means that ${\bf b}(i)\leq{\bf a}(i)$ for all $i$.
A monomial ideal $I\subset S$ generated in a single degree is said to satisfy the strong exchange property if: for all $u,v\in\mathcal{G}(I)$, all $i$ such that $\deg_{x_{i}}(u)>\deg_{x_{i}}(v)$ and all $j$ such that $\deg_{x_{j}}(u)<\deg_{x_{j}}(v)$ we have $x_{j}(u/x_{i})\in I$. It is clear that an ideal of Veronese type satisfies the strong exchange property. By the work [24], it is known that ideals satisfying the strong exchange property are essentially ideals of Veronese type in the sense that if $I$ satisfies the strong exchange property, then $I=(u)I_{{\bf a},d}$ for some monomial $u\in S$, and some ${\bf a}\in{\mathbb{Z}}_{\geq 0}^{n}$ and $d>0$.
Proposition 4.5.
Let $G$ be a tree with $n$ vertices. Then $J_{i}(G)$ is a poymatroidal ideal with the strong exchange property for all $i\in[n-2]$.
Proof.
Let $K_{i}(G)=\left(\prod_{k\in F}x_{\phi(k)}\mid F\subseteq[n-2]\text{\ and\ }|F|=i\right)$. For any $k\in[n-1]$, define
$b_{k}=|\phi^{-1}(k)|=\left|\{j\in[n-2]:\phi(j)=k\}\right|.$
Then $b_{k}>0$ if and only if $k\in\phi([n-2])$. Moreover, $\sum\limits_{i=1}^{n-1}b_{i}=n-2$. We claim that
$K_{i}(G)=\left(\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}\mid\sum\limits_{k=1}^{n-1}a_{k}=i\text{\ and\ }0\leq a_{k}\leq b_{k}\ \text{for all}\ k\in[n-1]\right).$
For every $F\subseteq[n-2]$ such that $|F|=i$, let $a_{k}=|F\cap\phi^{-1}(k)|$ for any $k\in[n-1]$. Then $0\leq a_{k}\leq b_{k}$ for all $k$, and
$\sum\limits_{k=1}^{n-1}a_{k}=\left|\bigsqcup_{k=1}^{n-1}(F\cap\phi^{-1}(k))\right|=|F|=i.$
Note that $\prod_{k\in F}x_{\phi(k)}=\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}$, so the inclusion $\subseteq$ holds.
Conversely, take any monomial $\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}$ such that $\sum\limits_{k=1}^{n-1}a_{k}=i$ and $0\leq a_{k}\leq b_{k}$ for all $k\in[n-1]$. Let $F=F_{1}\sqcup\cdots\sqcup F_{n-1}$, where $F_{k}$ is a subset of $a_{k}$ elements in $\phi^{-1}(k)$, and by convention, $F_{k}=\emptyset$ if $a_{k}=0$. Then $\prod_{k\in F}x_{\phi(k)}=\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}$, so the inclusion $\supseteq$ holds. This proves the claim.
Therefore, $K_{i}(G)$ is the ideal $I_{{\bf b};i}$ of Veronese type on the ground set $\phi([n-2])$. Here, ${\bf b}=(b_{1},\ldots,b_{n-1}).$ By putting $b_{n}=0$, we conclude that
$J_{i}(G)$ $=\left(\prod\limits_{k=1}^{n}x_{k}^{i-a_{k}}\mid\sum\limits_{k=1}^{n-1}a_{k}=i\text{\ and\ }0\leq a_{k}\leq b_{k}\ \text{for all}\ k\in[n]\right)$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{c_{k}}\mid\sum\limits_{k=1}^{n-1}c_{k}=(n-2)i\text{\ and\ }i-b_{k}\leq c_{k}\leq i\ \text{for all}\ k\in[n]\right)$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{i-b_{k}}\right)\left(\prod\limits_{k=1}^{n-1}x_{k}^{c_{k}-i+b_{k}}\mid\sum\limits_{k=1}^{n-1}c_{k}=(n-2)i\text{\ and\ }i-b_{k}\leq c_{k}\leq i\ \text{for all}\ k\in[n]\right)$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{i-b_{k}}\right)\left(\prod\limits_{k=1}^{n-1}x_{k}^{a_{k}}\mid\sum\limits_{k=1}^{n-1}a_{k}=n-2-i\text{\ and\ }0\leq a_{k}\leq b_{k}\ \text{for all}\ k\in[n]\right)$
$=\left(\prod\limits_{k=1}^{n}x_{k}^{i-b_{k}}\right)K_{n-2-i}(G).$
Here, the second to last equality holds because of $\sum\limits_{i=1}^{n-1}b_{i}=n-2$. In the last equality, we use the convention that $K_{0}(G)=S$. Therefore, $J_{i}(G)$ is a polymatroidal ideal with the strong exchange property.    ∎
In fact, $J_{i}(G)$ is the dual polymatoidal ideal of $K_{i}(G)$ with respect to the cage the $n$-tuple $(i,\ldots,i)$ in the following sense. Let $I$ be a polymatoidal ideal. A cage for $I$ is an integral vector ${\bf c}\in\mathbb{N}^{n}$ such that ${\bf u}(i)\leq{\bf c}(i)$ for any $\bf u$} with ${\bf x}^{\bf u}\in\mathcal{G}(I)$ and any $i\in[n]$. Fix a cage $\bf c$} of $I$. Then the ideal generated ${\bf x}^{\bf c-u}:\ {\bf u}\in G(I)$ is also a polymatroid ideal, which we call the dual polymatroidal ideal of $P$ with respect to $\bf c$}. For more details, see [6] and [13].
In Lemma 4.5, we prove that the $i$-th homological shift ideal $\mathrm{HS}_{i}(I_{c}(G)^{i})(=J_{i}(G))$, when dividing a suitable monomial, is a ideal of Veronese type. In fact that every Veronese-type ideal arises in this way. Recall that a caterpillar tree is a tree $T$ with the property that it contains a path $P$ such that any vertex of $T$ is either a vertex of $P$ or it is adjacent to a vertex of $P$. Caterpillar trees were first studied by Harary and Schwenk [21]. A leaf edge is an edge that connects a leaf vertex (a vertex of degree 1) to another vertex in the graph.
Theorem 4.6.
Let $I=I_{\mathbf{a},d}\subseteq\mathbb{K}[x_{1},\ldots,x_{n}]$ be a Veronese-type ideal, where $\mathbf{a}=(a_{1},\ldots,a_{n})\in\mathbb{Z}_{>0}^{n}$ and $d\leq|\mathbf{a}|$ (with $|\mathbf{a}|=a_{1}+\cdots+a_{n}$). Then there exists a caterpillar tree $T$ such that
$\mathrm{HS}_{|\mathbf{a}|-d}\left(I_{c}(T)^{|\mathbf{a}|-d}\right)=uI$
for some monomial $u$.
Proof.
Set $\sigma_{0}:=0$ and $\sigma_{k}:=\sum\limits_{j=1}^{k}a_{j}$ for all $k\in[n]$.
We now construct a caterpillar tree $T$ such that $K_{d}(T)=I$. Let $P$ be a path of length $n$ with $n+1$ vertices, labeled in order by $\sigma_{1}+1,\sigma_{2}+1,\ldots,\sigma_{n}+1,\sigma_{n}+2$. We then add leaf edges according to the sequence $(a_{1},\ldots,a_{n})$: attach $\sigma_{1}$ leaf edges to the vertex $\sigma_{1}+1$, whose leaf vertices are $1,\ldots,\sigma_{1}$; for each $2\leq i\leq n$, attach $\sigma_{i}-\sigma_{i-1}-1=a_{i}-1$ leaf edges to the vertex $\sigma_{i}+1$, with leaf vertices $\sigma_{i-1}+2,\ldots,\sigma_{i-1}+a_{i}=\sigma_{i}$.
Denote the resulting graph by $T$. By construction, $T$ is a caterpillar tree with vertex set $\sigma_{n}+2]$. Let $\phi:[\sigma_{n}+1]\rightarrow[\sigma_{n}+2]$ be the map where $\phi(k)$ is the unique vertex $j$ satisfying $j>k$ and $\{j,k\}\in E(T)$. For the role of $\phi$, see Theorem 4.3.
[…]
Figure 1. Caterpillar tree $T$
Then, for any $k\in[\sigma_{n}+2]$, the set $\phi^{-1}(k)$ is nonempty if and only if $k=\sigma_{j}+1$ for some $1\leq j\leq n$. Furthermore, for each $j\in[\sigma+1]$, we have $\phi^{-1}(\sigma_{j}+1)=\{\sigma_{j-1}+1,\sigma_{j-1}+2,\ldots,\sigma_{j}\}$. In particular, $|\phi^{-1}(\sigma_{j}+1)|=a_{j}$ for $j\in[n]$. By the proof of Proposition 4.5, $K_{d}(T)=I=I_{\mathbf{a},d}$. Moreover,
$\mathrm{HS}_{|\mathbf{a}|-d}(I_{c}(T)^{|\mathbf{a}|-d})=J_{|\mathbf{a}|-d}(T)=uK_{d}(T)=uI$
for some monomial $u$, as required.    ∎
4.3. Cycles
In this subsection, we investigate the homological shift algebras associated to the complementary edge ideal of a cycle.
Let $G$ be an $n$-cycle with vertices labeled $1,2,\ldots,n$ in cyclic order. For $j=1,\ldots,n-1$, let ${\bf e}_{j}$ denote the edge $x_{j}x_{j+1}$, and set ${\bf e}_{n}=x_{n}x_{1}$. For notational convenience, we also use ${\bf e}_{0}$ to denote ${\bf e}_{n}$ and $x_{0}$ to denote $x_{n}$.
We first consider the case when $G$ is an even cycle. Note that the maximal element (set) of the sets $\operatorname{set}(\beta)$, where $\beta\in\mathcal{G}(I^{s})$ for some $s\geq 1$ is equal to $\{1,2,\ldots,n-2\}$.
Lemma 4.7.
Let $G$ be the even cycle of size $n=2m$ and let $I$ be the complementary edge ideal of $G$. Take $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in I^{s}$, where $u_{j}\in I(G)$ for all $j$. Then
  1. (1)
    If $u_{1}={\bf e}_{2\ell-1}=x_{2\ell-1}x_{2\ell}$ for some $1\leq\ell\leq m-1$, then $2\ell$ belongs to $\operatorname{set}(\beta)$ if and only if either $u_{j}=x_{2\ell}x_{2\ell+1}$ for some $j\geq 2$ or $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-3}\}\subseteq\{u_{2},\ldots,u_{s}\}.$
  2. (2)
    If $u_{1}={\bf e}_{2\ell}=x_{2\ell}x_{2\ell+1}$ for some $1\leq\ell\leq m-2$, then $2\ell+1$ belongs to $\operatorname{set}(\beta)$ if and only if either $u_{j}=x_{2\ell+1}x_{2\ell+2}$ for some $j\geq 2$ or $\{{\bf e}_{0},{\bf e}_{2},{\bf e}_{4},\ldots,{\bf e}_{2\ell-2}\}\subseteq\{u_{2},\ldots,u_{s}\}.$
Proof.
(1) $\Rightarrow$ Suppose that $2\ell$ belongs to $\operatorname{set}(\beta)$ and $u_{j}\neq x_{2\ell}x_{2\ell+1}$ for any $2\leq j\leq s$. Then there exists $k\geq 2\ell$ such that $2\ell-1$ is even connected to $k$ with respect to $u_{2}\cdots u_{s}$ by the following walk:
$2\ell-1=y_{0}--y_{1}--y_{2}--\cdots--y_{2t-1}--y_{2t}--y_{2t+1}=k.$
Assume further that this walk cannot be shotten. We use induction on $\ell$ to show that $\{x_{1}x_{2},x_{3}x_{4},\ldots,x_{2\ell-3}x_{2\ell-2}\}\subseteq\{u_{2},\ldots,u_{s}\}.$ If $\ell=1$, there is nothing to prove, since $\{x_{1}x_{2},x_{3}x_{4},\ldots,x_{2\ell-3}x_{2\ell-2}\}$ is an empty set. Suppose now that $\ell>1$ It is clear that $y_{1}\in\{2\ell,2\ell-2\}$. But if $y_{1}=2\ell$, then $y_{2}$ has to be $2\ell-1$, for otherwise $u_{j}=y_{1}y_{2}=x_{2\ell}x_{2\ell+1}$ for some $j\geq 2$. This implies the above-mentioned walk can be shorten, a contradiction. Hence $y_{1}=2\ell-2$. From this, it follows that $y_{2}=2\ell-3$. Hence, $y_{1}y_{2}\in\{u_{2},\ldots,u_{s}\}$. We may assume $u_{2}=x_{2\ell-3}x_{2\ell-2}$ without loss of generality. Note that $2\ell-3=y_{2q+2}--y_{2q+3}--y_{2q+4}---\cdots--y_{2t-1}--y_{2t}--y_{2t+1}=k$ is a walk by which $2\ell-3$ is even connected to $k$ with respect to $u_{3}u_{4}\cdots u_{s}$. Note that $2\ell-1$ does not belongs to this walk. Hence, by induction, we conclude that $\{x_{1}x_{2},x_{3}x_{4},\ldots,x_{2\ell-5}x_{2\ell-4}\}\subseteq\{u_{3},\ldots,u_{s}\}$, completing the proof.
$\Leftarrow$ Assume that $u_{1}={\bf e}_{2\ell-1}=x_{2\ell-1}x_{2\ell}$. If $u_{j}=x_{2\ell}x_{2\ell+1}$ for some $2\leq j\leq s$, then, since $2\ell+2\leq 2m$ is a vertex of $G$ such that $2\ell+2$ is adjacent to $2\ell+1$ and $2\ell+2>2\ell$, we conclude that $2\ell\in\operatorname{set}(\beta)$ by Lemma 3.3. If $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-3}\}\subseteq\{u_{2},\ldots,u_{s}\},$ then $2m$ is even connected to $2\ell-1$ with respect to ${\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2\ell-3}$ by the following walk
$2m--\underbrace{1--2}-3--4--\cdots--\underbrace{(2\ell-3)--(2\ell-2)}--(2\ell-1).$
Hence, $2\ell$ belongs to $\operatorname{set}(\beta)$.
The proof of (2) is similar to the proof of (1), we omit it.    ∎
Next, we consider the case when $n$ is odd. Note that there exists $\beta\in\mathcal{G}(I^{s})$ for some $s\geq 1$ such that $\operatorname{set}(\beta)$ is equal to $\{1,2,\ldots,n-1\}$.
Lemma 4.8.
Let $G$ be the odd cycle of size $n=2m+1$ and let $I$ be the complementary edge ideal of $G$. Take $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in I^{s}$, where $u_{j}\in I(G)$ for all $j$. Then the following conclusions hold:
  1. (1)
    If $u_{1}={\bf e}_{2\ell-1}=x_{2\ell-1}x_{2\ell}$ for some $1\leq\ell\leq m-1$, then $2\ell$ belongs to $\operatorname{set}(\beta)$ if and only if either $u_{j}=x_{2\ell}x_{2\ell+1}$ for some $j\geq 2$ or $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-3}\}\subseteq\{u_{2},\ldots,u_{s}\}.$
  2. (2)
    If $u_{1}={\bf e}_{2\ell}=x_{2\ell}x_{2\ell+1}$ for some $1\leq\ell\leq m-2$, then $2\ell+1$ belongs to $\operatorname{set}(\beta)$ if and only if either $u_{j}=x_{2\ell+1}x_{2\ell+2}$ for some $j\geq 2$ or $\{{\bf e}_{n},{\bf e}_{2},{\bf e}_{4},\ldots,{\bf e}_{2\ell-2}\}\subseteq\{u_{2},\ldots,u_{s}\}.$
  3. (3)
    $2m\in\operatorname{set}(\beta)$ if and only if $\{{\bf e}_{1},{\bf e}_{3}\ldots,{\bf e}_{2m-3},{\bf e}_{2m-1}\}\subseteq\{u_{1},\ldots,u_{s}\}.$
Proof.
The proofs of (1) and (2) are similar to that of Lemma 4.7, so we only prove (3).
(3) Suppose that $\{{\bf e}_{1},{\bf e}_{3}\ldots,{\bf e}_{2m-3},{\bf e}_{2m-1}\}\subseteq\{u_{1},\ldots,u_{s}\}$. We may assume that $u_{1}={\bf e}_{2m-1}$ without loss of generality. Then, since $2m+1$ is even connected to 2m12m-1 with respect to $u_{2}\cdots u_{s}$ by the following walk:
$(2m+1)--1--2--3--4--\cdots--(2m-3)--(2m-2)--(2m-1),$
we conclude that $2m\in\operatorname{set}(\beta)$ by Lemma 3.3.
Conversely, if $2m\in\operatorname{set}(\beta)$, we may assume that $u_{1}\in\{x_{2m}x_{2m+1},x_{2m-1}x_{2m}\}$. We need to consider the following cases:
(i) $u_{1}=x_{2m}x_{2m+1}$: By Lemma 3.3, $2m+1$ is even connected to $2m+1$ with respect to $u_{2}u_{3}\cdots u_{s}$. We may assume the following walk is a shortest walk by which $2m+1$ even connected to $2m+1$ with respect to $u_{2}u_{3}\cdots u_{s}$:
$2m+1=y_{0}--y_{1}--y_{2}--\cdots--y_{2t-1}--y_{2t}--y_{2t+1}=2m+1,({\dagger})$
It is clear that $y_{1}\in\{1,2m\}$. We consider the case when $y_{1}=1$ only, since the proof of the other case is similar. Since $y_{1}=1$, we have $y_{2}\in\{2,2m+1\}$. But if $y_{2}=2m+1$, then $2m+1--y_{3}-y_{4}----\cdots--(2m-1)--(2m)--(2m+1)$ is also a walk by which $2m+1$ is even connected to $2m+1$. This is contradicted to the assumption that the walk ($\dagger$}) is shortest. Thus $y_{2}=2$. This implies $y_{3}=3$ or $y_{3}=1$. But if $y_{3}=1$, the walk ($\dagger$}) can be shortened, a contradiction again. It follows that $y_{3}=3$. By analogy, we conclude that $y_{4}=4,y_{5}=5,\cdots,$. Finally, we have $t=m$ and $y_{2m}=2m$. This implies that $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{2},u_{3},\ldots,u_{s}\}$. In particular, $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{1},u_{2},u_{3},\ldots,u_{s}\}$, as desired.
Remark: (1) In view of the proof of (i), we can conclude that $({\dagger})$ is actually the unique shortest walk through which vertex $2m+1$ is even connected to itself with respect to any product of edges. Moreover, any shortest walk that even connects a vertex in the cycle to itself should have the same structure.
(2) In the case when $G$ is an even cycle, we may prove that there are no walks by which a vertex is even connected to itself with respect to any product of edges, using a similar argument as above. This explains why $\max\{\mbox{pd}(I_{c}(G)^{s}):\ s\geq 1\}=|V(G)|-2$, rather than |V(G)|1|V(G)|-1.
(ii) $u_{1}=x_{2m-1}x_{2m}={\bf e}_{2m-1}$: By Lemma 3.3, $2m+1$ is even connected to 2m12m-1 with respect to $u_{2}u_{3}\cdots u_{s}$. We may assume the following walk is a shortest walk by which $2m+1$ is even connected to 2m12m-1 with respect to $u_{2}u_{3}\cdots u_{s}$:
$2m+1=y_{0}--y_{1}--y_{2}--\cdots--y_{2t-1}--y_{2t}--y_{2t+1}=2m-1.$
If $y_{1}=2m$, then $y_{2}=2m-1$. It follows that $2m-1--y_{3}--y_{4}--y_{5}--y_{6}--\cdots---y_{2t-1}--y_{2t}--y_{2t+1}=2m-1$ is a shortest walk that even connects 2m12m-1 to 2m12m-1 with respect to $u_{2}\cdots u_{2}$. This walk must have the same structure as the walk that even connects $2m+1$ to $2m+1$ discussed in Case (i). Hence, it is $(2m-1)--(2m)--(2m+1)--1--2--3--4--\cdots--(2m-3)--(2m-2)--(2m-1)$, and in particular, $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{2},u_{3},\ldots,u_{s}\}$.
It remains to consider the case when $y_{1}=1$. In this case, we can conclude, as in case (i), that $y_{2}=2,y_{3}=3,\cdots,$, and that t=m1t=m-1, $y_{2m}=2m-2$. Hence, we have $\{{\bf e}_{1},{\bf e}_{3},\cdots{\bf e}_{2m-1}\}\subseteq\{u_{2},u_{3},\ldots,u_{s}\}$ again.    ∎
Lemma 4.9.
Let $G$ be the cycle of size $n$, whose vertices are labelled $1,2,\ldots,n$ successively. Let ${\bf e}_{j}$ denote the edge $x_{j}x_{j+1}$ for $j=1,\ldots,n-1$ and set ${\bf e}_{0}={\bf e}_{n}=x_{n}x_{1}$. Take $\beta=\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}\in I^{s}$. Let $T=\{u_{1},\ldots,u_{s}\}$. Define
$N_{1}(\beta)=\left\{\begin{array}[]{ll}\max\{[2\ell]:\ \{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-1}\}\subseteq T,1\leq\ell\leq m-1\},&\hbox{ $n=2m$;}\\ \max\{[2\ell]:\ \{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-1}\}\subseteq T,1\leq\ell\leq m\},&\hbox{ $n=2m+1$.}\end{array}\right$.
$N_{2}(\beta)=\left\{\begin{array}[]{ll}\max\{[2\ell+1]:\ \{{\bf e}_{0},{\bf e}_{2},\ldots,{\bf e}_{2\ell}\}\subseteq T,\ 1\leq\ell\leq m-2\},&\hbox{$n=2m$;}\\[2.0pt] \max\{[2\ell+1]:\ \{{\bf e}_{0},{\bf e}_{2},\ldots,{\bf e}_{2\ell}\}\subseteq T,\ 1\leq\ell\leq m-1\},&\hbox{$n=2m+1$.}\end{array}\right$.
By convention, if ${\bf e}_{1}\notin\{u_{1},\ldots,u_{s}\}$ then $N_{1}(\beta)=\emptyset$ and if ${\bf e}_{0}\notin\{u_{1},\ldots,u_{s}\}$ then $N_{2}(\beta)=\emptyset$. Then
$\operatorname{set}(\beta)=N_{1}(\beta)\cup N_{2}(\beta)\cup\left(\{\min{u_{j}}:\ 1\leq j\leq s\}\setminus\{n-1\}\right).$
Before the proof, we made some remarks. If $n=2m+1$ is an odd number, any $\beta\in\mathcal{G}(I^{s})$ is expressed as $\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}$ uniquely. Hence, there is no ambiguity in the definitions of $N_{1}(\beta)$ and $N_{2}(\beta)$. However, if $n$ is even, we should note that the elements of $N_{1}(\beta)$ and $N_{2}(\beta)$ depend on the way $\beta$ is written as $\frac{\alpha^{s}}{u_{1}u_{2}\cdots u_{s}}$. For example, let $\beta_{1}=\frac{\alpha^{m}}{{\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2m-1}}$ and $\beta_{2}=\frac{\alpha^{m}}{{\bf e}_{2}{\bf e}_{4}\cdots{\bf e}_{2m}}$. Then $\beta_{1}=\beta_{2}$. However, $N_{1}(\beta_{1})=[2m-2]$, $N_{1}(\beta_{2})=\emptyset$, $N_{2}(\beta_{1})=\emptyset$ and $N_{2}(\beta_{2})=[2m-3]$. Of course, $\operatorname{set}(\beta)$ is independent of the expression of $\beta$. In the above example, note that $\min{\bf e}_{2m-2}=2m-2$. we have $\operatorname{set}(\beta_{1})=\operatorname{set}(\beta_{2})=[2m-2]$. Since the only real relation that occurs among ${\bf e}_{1},\ldots,{\bf e}_{n}$ is ${\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2m-1}={\bf e}_{2}\cdots{\bf e}_{2m-2}{\bf e}_{2m}$, the set $\operatorname{set}(\beta)$ given in Lemma 4.9 is actually independent of the expression of $\beta$.
Proof.
We only consider the case when $n=2m+1$ (an odd number), as the proof for the other case is analogous and even simpler. Let $j\in\mathrm{set}(\beta)$. Then $j\in[n-1]$ and $j\in u_{k}$ for some $1\leq k\leq s$. If j=n1=2mj=n-1=2m, then by Lemma 4.8.(3), $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2m-3},{\bf e}_{2m-1}\}\subseteq\{u_{1},\ldots,u_{s}\}$, which implies $j\in N_{1}(\beta)$. Thus, we may assume $j\leq 2m-1$.
If $j=\min u_{k}$ for some $k\in[s]$, then $j\in\{\min u_{j}:\ 1\leq j\leq s\}\setminus\{n-1\}$. Suppose now that $j\neq\min u_{k}$ for any $k=1,\ldots,s$. By Lemma 4.8, we conclude:
- If $j=2\ell$ for some $\ell$, then $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-1}\}\subseteq\{u_{1},\ldots,u_{s}\}$;
- If $j=2\ell+1$ for some $\ell$, then $\{{\bf e}_{n},{\bf e}_{2},\ldots,{\bf e}_{2\ell}\}\subseteq\{u_{1},\ldots,u_{s}\}$.
Since $j\leq 2m-1$, we have: If $j=2\ell+1$, then $\ell\leq m-1$; If $j=2\ell$, then $\ell\leq m-1$. This implies $j\in N_{1}(\beta)\cup N_{2}(\beta)$, so the inclusion $\subseteq$ holds. The converse inclusion follows directly from Lemma 4.8.
As an immediate consequence of Lemma 4.9, we obtain:
Corollary 4.10.
Keep the notation as before. Let $\beta=\frac{\alpha^{s}}{{\bf e}_{1}^{s_{1}}{\bf e}_{2}^{s_{2}}\cdots{\bf e}_{n}^{s_{n}}}\in\mathcal{G}(I^{s})$. Define $t_{j}=1$ if $s_{j}>0$ and $t_{j}=0$ if $s_{j}=0$, for $1\leq j\leq n$. Then $\operatorname{set}(\beta)=\operatorname{set}(\gamma)$, where $\gamma=\frac{\alpha^{t}}{{\bf e}_{1}^{t_{1}}{\bf e}_{2}^{t_{2}}\cdots{\bf e}_{n}^{t_{n}}}$ and $t=t_{1}+\cdots+t_{n}$.
By the above corollary, it is easy to see that $\mathrm{HS}_{i}(\mathcal{R(}I))$ is generated in degree $\leq n$ for all $i\geq 1$.
Hereafter, we devote to compute the homological shift ideals of powers of the complementary edge ideal of a cycle.
Lemma 4.11.
Let $G$ be a cycle of size $n$ and $I$ be its complementary edge ideal. For $1\leq i\leq n-1$ and $s\geq\frac{i}{2}$, the following inclusion holds:
(4) $\mathrm{HS}_{i}(I^{s})\subseteq\sum_{0\leq k\leq\frac{i}{2}}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\right).$
Here, $I^{t}=0$ if $t<0$.
Proof.
We only prove the case where $n=2m$ (even cycle), as the argument for $n=2m+1$ (odd cycle) is analogous. Fix $i\in[2m-2]$ and $s\geq\frac{i}{2}$. Let $\beta\in\mathcal{G}(I^{s})$ with $i\leq|\mathrm{set}(\beta)|$, and take a subset $Q\subseteq\mathrm{set}(\beta)$ with $|Q|=i$. It suffices to show that ${\bf x}_{Q}\beta$ lies in the right-hand side of (4).
Write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$ with $u_{j}\in I(G)$ for all $j=1,\ldots,s$. Let $T=\{u_{1},\ldots,u_{s}\}$. Denote
$\ell_{1}:=\max\left\{1\leq\ell\leq m-1:\ \{{\bf e}_{1},{\bf e}_{3},\ldots{\bf e}_{2\ell-1}\}\subseteq T\right\}$
and
$\ell_{2}:=\max\{1\leq\ell\leq m-2:\ \{{\bf e}_{0},{\bf e}_{2},\ldots{\bf e}_{2\ell}\}\subseteq T\}.$
By convention, $\max\emptyset=$-$\infty$. It is clear that $N_{1}(\beta)\cup N_{2}(\beta)=[2\ell_{1}]\cup[2\ell_{2}+1]$. We consider two cases:
(1) $Q\cap\left([2\ell_{1}]\cup[2\ell_{2}+1]\right)=\emptyset$: By Lemma 4.9, there exist distinct edges $v_{1},\ldots,v_{i}\in T$ such that $Q=\{\min(v_{1}),\ldots,\min(v_{i})\}$. We may write $u_{1}u_{2}\cdots u_{s}=(v_{1}\cdots v_{i})(u_{1}\cdots u_{s-i})$ (after reindexing if necessary). Note that $\max(v_{1}),\ldots,\max(v_{i})$ are pairwise distinct; let $F$ denote the $i$-element set $\{\max(v_{1}),\ldots,\max(v_{i})\}$.
Then
${\bf x}_{Q}\beta=\frac{\alpha^{s}}{{\bf x}_{F}\cdot u_{1}\cdots u_{s-i}}=\frac{\alpha^{i}}{{\bf x}_{F}}\cdot\frac{\alpha^{s-i}}{u_{1}\cdots u_{s-i}}.$
Hence, ${\bf x}_{Q}\beta$ belongs to the $k=0$ component of the right-hand side of (4).
(2) $Q\cap\left([2\ell_{1}]\cup[2\ell_{2}+1]\right)=Q_{1}\neq\emptyset$: In this case, we can decompose $Q$ as
$Q=Q_{1}\sqcup\{j_{p+1}<\cdots<j_{i}\},$
where $|Q_{1}|=p$, $j_{p+1}>\max Q_{1}$, and for all $k=p+1,\ldots,i$,
${\bf e}_{j_{k}}\in\{u_{1},u_{2},\ldots,u_{s}\}\setminus\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell_{1}-1},{\bf e}_{0},{\bf e}_{2},\ldots,{\bf e}_{2\ell_{2}}\}.$
Note that if $p=i$, then $\{j_{p+1}<\cdots<j_{i}\}$ is an empty set.
First, consider the subcase $\ell_{2}\geq\ell_{1}$. Here, $Q_{1}\subseteq{\bf e}_{0}\cup{\bf e}_{2}\cup\cdots\cup{\bf e}_{2\ell_{2}}$. Let $\Omega:=\{{\bf e}_{0},\ldots,{\bf e}_{2\ell_{2}}\}$. There exist edges $v_{1},\ldots,v_{q}\in\Omega\setminus\{{\bf e}_{0}\}$ and edges $f_{1},\ldots,f_{r}\in\Omega$ such that
${\bf x}_{Q_{1}}=v_{1}v_{2}\cdots v_{q}\cdot x_{j_{1}}\cdots x_{j_{r}},$
where $j_{t}\in f_{t}$ for $t=1,\ldots,r$. These edges $v_{1},\ldots,v_{q},f_{1},\ldots,f_{r}$ are pairwise disjoint, with $2q+r=p$ and $q+r\leq\ell_{2}$. For each $t=1,\ldots,r$, let $i_{t}\in[2\ell_{2}+1]$ be such that $f_{t}=x_{j_{t}}x_{i_{t}}$.
Since $\{v_{1},\ldots,v_{q},f_{1},\ldots,f_{r},{\bf e}_{j_{p+1}},\ldots,{\bf e}_{j_{i}}\}\subseteq\{u_{1},\ldots,u_{s}\}$, we may assume (after reindexing) that
$u_{1}\cdots u_{s}=\left(v_{1}\cdots v_{q}f_{1}\cdots f_{r}{\bf e}_{j_{p+1}}\cdots{\bf e}_{j_{i}}\right)u_{1}\cdots u_{s-i+q}.$
Then
${\bf x}_{Q}\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s-i+q}\cdot(x_{i_{1}}\cdots x_{i_{r}})\cdot(x_{j_{p+1}+1}\cdots x_{j_{i}+1})}.$
Let $F=\{i_{1},\ldots,i_{r},j_{p+1}+1,\ldots,j_{i}+1\}$, which is an (iq)(i-q)-subset of $[n]$. Thus, ${\bf x}_{Q}\beta\in\frac{\alpha^{i-q}}{{\bf x}_{F}}I^{s-i+q}$, meaning ${\bf x}_{Q}\beta$ belongs to the $k=q$ component of the right-hand side of (4).
The proof for $\ell_{1}>\ell_{2}$ is analogous, so the argument is complete.    ∎
Next, we show that if $i$ is large enough, i.e., $i>\frac{n}{2}$, certain components of the right-hand of (4) may be removed.
Proposition 4.12.
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. For all integers $1\leq i<n$ and $s\geq\left\lfloor\frac{i}{2}\right\rfloor$, the following inclusion holds:
(5) $\mathrm{HS}_{i}(I^{s})\subseteq\sum_{\begin{subarray}{c}\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor\end{subarray}}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\right).$
Here, $I^{t}=0$ for any integer $t<0$ (by convention).
Proof.
If $i\leq\left\lceil\frac{n}{2}\right\rceil-1$, then the right-hand side of (5) coincides with the left-hand side of (4). Thus, the result follows from Lemma 4.11. Suppose now that $i>\left\lceil\frac{n}{2}\right\rceil-1$. We only consider the case where $n=2m$, as the argument for $n=2m+1$ is analogous. We may write i=m1+qi=m-1+q. Note that $q=\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}$. Let $\beta\in\mathcal{G}(I^{s})$ with $i\leq|\mathrm{set}(\beta)|$, and take a subset $Q\subseteq\mathrm{set}(\beta)$ with $|Q|=i$. It suffices to show that ${\bf x}_{Q}\beta$ lies in the right-hand side of (5). Write $\beta=\frac{\alpha^{s}}{u_{1}\cdots u_{s}}$ with $u_{j}\in I(G)$ for all $j=1,\ldots,s$. Let $T=\{u_{1},\ldots,u_{s}\}$.
Denote
$\ell_{1}:=\max\left\{1\leq\ell\leq m-1:\ \{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell-1}\}\subseteq T\right\}$
and
$\ell_{2}:=\max\left\{1\leq\ell\leq m-2:\ \{{\bf e}_{0},{\bf e}_{2},\ldots,{\bf e}_{2\ell}\}\subseteq T\right\}.$
By convention, $\max\emptyset=$-$\infty$. Suppose first that $\ell_{1}>\ell_{2}$. By Lemma 4.9, it follows that
$\mathrm{set}(\beta)=[2\ell_{1}]\cup\left(\{\min u_{j}:\ j=1,\ldots,s\}\setminus\{n-1\}\right).$
Note that the edges ${\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2m-3}$ are pairwise disjoint, and their union is the set [2m2][2m-2]. Since $Q\subseteq\mathrm{set}(\beta)\subseteq[2m]$ and |Q|=m1+q|Q|=m-1+q, by the pigeonhole principle, there are $k\geq q$ edges among ${\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2m-3}$ that are entirely contained in $Q$. Let these $k$ edges be ${\bf e}_{2j_{1}-1},{\bf e}_{2j_{2}-1},\ldots,{\bf e}_{2j_{k}-1}$ with $j_{1}<j_{2}<\cdots<j_{k}$.
There are three possible scenarios for the indices $j_{t}$ relative to $\ell_{1}$: 1. $\ell_{1}<j_{1}$, 2. $j_{k}\leq\ell_{1}$, 3. There exists some $1\leq r\leq k-1$ such that $j_{r}\leq\ell_{1}<j_{r+1}$ (indices straddle $\ell_{1}$).
We only prove the third scenario, as the first two follow by analogous reasoning. For $r+1\leq t\leq k$, since $j_{t}>\ell_{1}$, we have $2j_{t}-1>2\ell_{1}$. Given that $2j_{t}-1\in Q\subseteq\mathrm{set}(\beta)$, Lemma 4.9 implies ${\bf e}_{2j_{t}-1}\in T$. For $1\leq t\leq r$, since $j_{t}\leq\ell_{1}$, the edge ${\bf e}_{2j_{t}-1}$ is in $\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell_{1}-1}\}$, which is contained in $T$ by the definition of $\ell_{1}$. Thus, ${\bf e}_{2j_{1}-1},\ldots,{\bf e}_{2j_{k}-1}$ are pairwise disjoint edges in $T$ whose union is contained in $Q$. Let $Q_{1}=Q\setminus({\bf e}_{2j_{1}-1}\sqcup\cdots\sqcup{\bf e}_{2j_{k}-1})$. Then $|Q_{1}|=i-2k$; write $Q_{1}=\{\theta_{1},\ldots,\theta_{p}\}$ where p=i2kp=i-2k. For each $\theta_{j}\in Q_{1}$:
- If $\theta_{j}\leq 2\ell_{1}$, choose $v_{j}\in\{{\bf e}_{1},{\bf e}_{3},\ldots,{\bf e}_{2\ell_{1}-1}\}$ such that $\theta_{j}\in v_{j}$;
- If $\theta_{j}>2\ell_{1}$, set $v_{j}=\{\theta_{j},\theta_{j}+1\}$.
It is straightforward to verify that each $v_{j}\in T$. Let $F=\bigsqcup_{j=1}^{p}\left(v_{j}\setminus\{\theta_{j}\}\right)$. After reindexing, we may write
$u_{1}\cdots u_{s}=\left({\bf e}_{2j_{1}-1}\cdots{\bf e}_{2j_{k}-1}\right)\left(v_{1}\cdots v_{p}\right)\left(u_{1}\cdots u_{s-i+k}\right).$
Then
${\bf x}_{Q}\beta=\frac{\alpha^{s}}{(u_{1}\cdots u_{s-i+k})\cdot{\bf x}_{F}}\in\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}.$
Hence, ${\bf x}_{Q}\beta$ belongs to the right-hand side of (5).
Next, we consider the case when $\ell_{1}\leq\ell_{2}$, which is much more complicated than the first case. By Lemma 4.9, we have
$\mathrm{set}(\beta)=[2\ell_{2}+1]\cup\left(\{\min u_{j}:\ j=1,\ldots,s\}\setminus\{n-1\}\right).$
Consider the m2m-2 edges ${\bf e}_{2},{\bf e}_{4},\ldots,{\bf e}_{2m-4}$. Note that $\mathrm{set}(\beta)\subseteq({\bf e}_{2}\sqcup{\bf e}_{4}\sqcup\cdots\sqcup{\bf e}_{2m-4})\cup\{1,2m-2\}$, and since $Q\subseteq\mathrm{set}(\beta)$, it follows that
$Q\subseteq({\bf e}_{2}\sqcup{\bf e}_{4}\sqcup\cdots\sqcup{\bf e}_{2m-4})\cup\{1,2m-2\}.$
This inclusion implies
$|Q\cap({\bf e}_{2}\sqcup{\bf e}_{4}\sqcup\cdots\sqcup{\bf e}_{2m-4})|\geq|Q|-2.$
Since |Q|=m1+q|Q|=m-1+q, substituting this in gives
$|Q\cap({\bf e}_{2}\sqcup{\bf e}_{4}\sqcup\cdots\sqcup{\bf e}_{2m-4})|\geq(m-1+q)-2=m-3+q.$
Equality holds if and only if $Q\cap\{1,2m-2\}=\{1,2m-2\}$, i.e., $\{1,2m-2\}\subseteq Q$.
If $|Q\cap({\bf e}_{2}\sqcup{\bf e}_{4}\sqcup\cdots\sqcup{\bf e}_{2m-4})|\geq m-2+q$, then by the pigeonhole principle, there are at least $q$ edges among ${\bf e}_{2},{\bf e}_{4},\ldots,{\bf e}_{2m-4}$ that are entirely contained in $Q$. The proof then proceeds analogously to the case where $\ell_{1}>\ell_{2}$.
Assume now that $|Q\cap({\bf e}_{2}\sqcup{\bf e}_{4}\sqcup\cdots\sqcup{\bf e}_{2m-4})|=m-3+q$, which implies $\{1,2m-2\}\subseteq Q$. By the pigeonhole principle, there are at least q1q-1 edges among ${\bf e}_{2},{\bf e}_{4},\ldots,{\bf e}_{2m-4}$ that are entirely contained in $Q$. If there are more than q1q-1 such edges, we are done (as we would have at least $q$ edges). Thus, we may assume there are exactly q1q-1 such edges, denoted ${\bf e}_{2j_{1}},\ldots,{\bf e}_{2j_{q-1}}$ where $j_{1}<j_{2}<\cdots<j_{q-1}$. For the remaining edges ${\bf e}_{2t}$ with $t\in[m-2]\setminus\{j_{1},\ldots,j_{q-1}\}$, it follows that $|Q\cap{\bf e}_{2t}|=1$.
The indices $j_{1},\ldots,j_{q-1}$ satisfy one of three scenarios relative to $\ell_{2}$: 1. $\ell_{2}<j_{1}$ (all indices are greater than $\ell_{2}$), 2. $j_{q-1}\leq\ell_{2}$ (all indices are less than or equal to $\ell_{2}$), 3. There exists some $1\leq r\leq q-2$ such that $j_{r}\leq\ell_{2}<j_{r+1}$ (indices straddle $\ell_{2}$).
We focus on the third scenario (the others follow similarly, with minor adjustments). First, we claim $2\ell_{2}+2\notin Q$: if $2\ell_{2}+2\in Q$, then $2\ell_{2}+2\in\mathrm{set}(\beta)$, so Lemma 4.9 implies ${\bf e}_{2\ell_{2}+2}\in T$, contradicting the maximality of $\ell_{2}$.
Set p=q1rp=q-1-r. Since $2m-2\in Q$, we calculate the size of $Q\cap[2\ell_{2}+3,2m-2]$:
$|Q\cap[2\ell_{2}+3,2m-2]|=(m-1+q)-|Q\cap[1,2\ell_{2}+2]|.$
By construction, $|Q\cap[1,2\ell_{2}+2]|=|Q\cap[1,2\ell_{2}+1]|=\ell_{2}+r+1$ (including $1$, plus $2r$ elements from the $r$ edges ${\bf e}_{2j_{1}},\ldots,{\bf e}_{2j_{r}}\subseteq[2\ell_{2}+2]$, and $\ell_{2}-r$ elements from $\ell_{2}-r$ edges ${\bf e}_{2t}$ with $1\leq t\leq\ell_{2}$ and $t\notin\{j_{1},\ldots,j_{r}\}$). Substituting this gives:
$|Q\cap[2\ell_{2}+3,2m-2]|=(m-1+q)-(\ell_{2}+r+1)=m-\ell_{2}-1+p$
(using p=q1rp=q-1-r and simplifying).
Next, consider the $m-2-\ell_{2}$ edges ${\bf e}_{2\ell_{2}+3},{\bf e}_{2\ell_{2}+5},\ldots,{\bf e}_{2m-3}$, whose disjoint union is exactly $2\ell_{2}+3,2m-2]$. Since $|Q\cap[2\ell_{2}+3,2m-2]|=(m-2-\ell_{2})+(p+1)$, the pigeonhole principle guarantees at least $p+1$ of these edges are fully contained in $Q$. These $p+1$ edges, together with the $r$ edges ${\bf e}_{2j_{1}},\ldots,{\bf e}_{2j_{r}}$, form $(p+1)+r=q$ pairwise disjoint edges in $T$ whose union is contained in $Q$. We denote such $q$ edges by $f_{1},\ldots,f_{q}$. Let $Q_{1}$ be the remainder of $Q$ after removing all vertices contained in these $q$ edges.
Write $Q_{1}=\{\theta_{1},\ldots,\theta_{p}\}$ where p=i2qp=i-2q. For each $\theta_{j}\in Q_{1}$:
- If $\theta_{j}\leq 2\ell_{1}$, choose $v_{j}\in\{{\bf e}_{0},{\bf e}_{2},\ldots,{\bf e}_{2\ell_{2}}\}$ such that $\theta_{j}\in v_{j}$;
- If $\theta_{j}>2\ell_{2}+1$, set $v_{j}=\{\theta_{j},\theta_{j}+1\}$.
It is straightforward to verify that each $v_{j}\in T$. Let $F=\bigsqcup_{j=1}^{p}\left(v_{j}\setminus\{\theta_{j}\}\right)$. Then |F|=p=i2q|F|=p=i-2q. After reindexing, we may write
$u_{1}\cdots u_{s}=\left(f_{1}\cdots f_{q}\right)\left(v_{1}\cdots v_{p}\right)\left(u_{1}\cdots u_{s-i+q}\right).$
Then
${\bf x}_{Q}\beta=\frac{\alpha^{s}}{(u_{1}\cdots u_{s-i+q})\cdot{\bf x}_{F}}\in\frac{\alpha^{i-q}}{{\bf x}_{F}}\cdot I^{s-i+q}.$
Hence, ${\bf x}_{Q}\beta$ belongs to $k=q$ component of the right-hand side of (5). This completes the proof.    ∎
Let us consider the converse inclusion.
Lemma 4.13.
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. Assume $n\in\{2m,2m+1\}$. Given a subset $F\subseteq[n]$ with $p=|F|$ and an integer $k\geq 0$ such that $k+p<\frac{n}{2}$ (i.e., $k+p\leq m-1$ if $n=2m$, and $k+p\leq m$ if $n=2m+1$), set $i=2k+p$. Then there exist pairwise distinct edges $u_{1},\ldots,u_{k},u_{k+1},\ldots,u_{k+p}$ and a subset $E\subseteq\mathrm{set}\left(\frac{\alpha^{k+p}}{u_{1}\cdots u_{k+p}}\right)$ with $|E|=i$ such that $\frac{\alpha^{k+p}}{{\bf x}_{F}}={\bf x}_{E}\cdot\frac{\alpha^{k+p}}{u_{1}\cdots u_{k+p}}\in\mathrm{HS}_{i}(I^{k+p})$.
Proof.
We only consider the case $n=2m$; the proof for $n=2m+1$ is analogous. Let $F=\{j_{1}<j_{2}<\cdots<j_{p}\}$. We analyze two cases:
(1) $n\notin F$ (i.e., $j_{p}<n$): Define $t=\max\{a:j_{1}\leq 2k+2,\ j_{2}\leq 2k+4,\ \ldots,\ j_{a}\leq 2k+2a\}$. By convention, $t=0$ if $j_{1}>2k+2$. Then $1\leq j_{1}<\cdots<j_{t}\leq 2k+2t\leq 2m-2$ and $j_{p}>\cdots>j_{t+1}>2k+2t+2$. Set
$\beta=\frac{\alpha^{k+p}}{({\bf e}_{1}{\bf e}_{3}\cdots{\bf e}_{2k+2t-1})({\bf e}_{j_{t+1}-1}\cdots{\bf e}_{j_{p}-1})}.$
Then $E=([2k+2t]\setminus\{j_{1},\ldots,j_{t}\})\cup\{j_{t+1}-1,\ldots,j_{p}-1\}$ is a subset of $\mathrm{set}(\beta)$. Moreover, $|E|=2k+p=i$ and ${\bf x}_{E}\cdot\beta=\frac{\alpha^{k+p}}{{\bf x}_{F}}$, as required.
(2) $n\in F$ (i.e., $j_{p}=n$): As in case (1), there exists $t$ with $0\leq t\leq p-1$ such that $\{j_{1},\ldots,j_{t}\}\subseteq[2k+2t+1]\subseteq[2m-1]$ but $j_{t+1}>2k+2t+3$. Define
$\beta=\frac{\alpha^{k+p}}{({\bf e}_{0}{\bf e}_{2}\cdots{\bf e}_{2k+2t})({\bf e}_{j_{t+1}-1}\cdots{\bf e}_{j_{p-1}-1})}$
and
$E=\left([2k+2t+1]\setminus\{j_{1},\ldots,j_{t}\}\right)\cup\{j_{t+1}-1,\ldots,j_{p-1}-1\}.$
Then $|E|=2k+p$ and $E\subseteq\mathrm{set}(\beta)$. Furthermore, $\frac{\alpha^{k+p}}{{\bf x}_{F}}={\bf x}_{E}\cdot\beta$, as required.    ∎
We now are ready to present the main result of this subsection.
Theorem 4.14.
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. Then, for all integers $1\leq i<n$ and $s\geq\left\lfloor\frac{i}{2}\right\rfloor$ , the following equality holds:
(6) $\mathrm{HS}_{i}(I^{s})=\sum_{\begin{subarray}{c}\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor\end{subarray}}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\right)$
Proof.
For any $k$ satisfying
$\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor,$
let p=i2kp=i-2k. Then $k+p\leq\left\lceil\frac{n}{2}\right\rceil-1$ and k+p=ikk+p=i-k. By Lemma 4.13, for any $F\subseteq[n]$ with $|F|=p$, we have $\frac{\alpha^{i-k}}{{\bf x}_{F}}\in\mathrm{HS}_{i}(I^{i-k})$. It follows from [18, Proposition 1.3] that
$\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\subseteq\mathrm{HS}_{i}(I^{s}).$
Thus, all terms in the right-hand side are contained in $\mathrm{HS}_{i}(I^{s})$, so $\supseteq$ holds. The converse inclusion follows from Proposition 4.11. This completes the proof.    ∎
As an immediate consequence of Theorem 6, we obtain the generating degree of $\operatorname{HS}_{i}(\mathcal{R}(I))$.
Corollary 4.15.
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal. For all integers $1\leq i<\frac{n}{2}$, the $i$-th homological shift algebra $\operatorname{HS}_{i}(\mathcal{R}(I))$ is generated in degree $\leq i-q$. Here, $q=\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,0\right\}$.
Let $L_{j}$ denote the path with vertex set $[n]$ and edge set $E(G)\setminus\{j-1,j\}$ for $j=1,\ldots,n$. As before, 0 refers to the vertex $n$. Note that $L_{1},\ldots,L_{n}$ are all the spanning trees of $G$.
Let $I_{j}$ denote the complementary edge ideal of $L_{j}$ for all $j$. By Lemma 1.3, we have $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})\subseteq\mathrm{HS}_{i}(I^{s})$. It is therefore natural to ask which part of the right-hand of Formula (5) corresponds to $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})$.
Remark 4.16.
If $i<\frac{n}{2}$, then
$\mathrm{HS}_{i}(I^{s})=\sum_{0\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor}\left(\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i-2k\end{subarray}}\frac{\alpha^{i-k}}{{\bf x}_{F}}\cdot I^{s-i+k}\right).$
We claim that $\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})$ corresponds to the $k=0$ component of the expression above. More precisely,
$\sum_{j=1}^{n}\mathrm{HS}_{i}(I_{j}^{s})=\sum_{\begin{subarray}{c}F\subseteq[n]\\ |F|=i\end{subarray}}\frac{\alpha^{i}}{{\bf x}_{F}}\cdot I^{s-i}\quad\text{for all }s\geq i.$
In fact, by Theorem 4.3,
$\mathrm{HS}_{i}(I_{1}^{s})=\sum_{\begin{subarray}{c}F\subseteq[n]\setminus\{1,n\}\\ |F|=i\end{subarray}}\frac{\alpha^{i}}{\mathbf{x}_{F}}I^{s-i},$
and for $j=2,\ldots,n$,
$\mathrm{HS}_{i}(I_{j}^{s})=\sum_{\begin{subarray}{c}F\subseteq[n]\setminus\{j-1,j\}\\ |F|=i\end{subarray}}\frac{\alpha^{i}}{\mathbf{x}_{F}}I^{s-i}.$
This establishes the inclusion $\subseteq$ in the desired equality. For the converse, take any $F\subseteq[n]$ with $|F|=i$. Since every vertex cover of $G$ contains at least $m$ vertices (and $i<\frac{n}{2}=m$ for even $n$, or $i<\frac{n-1}{2}<m$ for odd $n$), $F$ cannot be a vertex cover of $G$. Thus, there exists some $j_{0}\in[n]$ such that $F\cap\{j_{0}-1,j_{0}\}=\emptyset$, which implies $\frac{\alpha^{i}}{{\bf x}_{F}}\cdot I^{s-i}\subseteq\mathrm{HS}_{i}(I_{j_{0}}^{s})$. This proves the claim.
For the largest $i$ satisfying $\mathrm{HS}_{i}(\mathcal{R}(I))\neq 0$, $\mathrm{HS}_{i}(I^{s})$ admits a simple expression.
Proposition 4.17.
Let $G$ be a cycle of length $n$ and $I$ be its complementary edge ideal.
(1) If $n=2m+1$, then $\mathrm{HS}_{2m}(I^{s})=\alpha^{m}\cdot I^{s-m}$ for $s\geq m$. In particular, $\mathrm{HS}_{2m}(\mathcal{R}(I))$ is generated in degree $m$.
(2) If $n=2m$, then $\mathrm{HS}_{2m-2}(I^{s})=\alpha^{m-1}\cdot I^{s-m+1}$ for $s\geq m-1$. In particular, $\mathrm{HS}_{2m-2}(\mathcal{R}(I))$ is generated in degree m1m-1.
Proof.
(1) In Formula (5), the range of $k$ is given by
$\max\left\{i-\left\lceil\frac{n}{2}\right\rceil+1,\ 0\right\}\leq k\leq\left\lfloor\frac{i}{2}\right\rfloor.$
For $n=2m+1$ and $i=2m$, substituting these values gives $\left\lceil\frac{n}{2}\right\rceil=m+1$, so the left bound simplifies to $\max\{2m-(m+1)+1,\ 0\}=\max\{m,\ 0\}=m$. The right bound is $\left\lfloor\frac{2m}{2}\right\rfloor=m$, hence $k=m$ is the only possible value. Thus, the result follows from Formula (5).
The proof of (2) is analogous.    ∎
Acknowledgments
The first author is supported by the National Natural Science Foundation of China (No. 11971338). The third author is supported by the Natural Science Foundation of Jiangsu Province (No. BK20221353) and the National Natural Science Foundation of China (No. 12471246).
Data availability statement
The data used to support the findings of this study are included within the article.
Conflict of interest
The authors declare that they have no competing interests.